### v1

# CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic

Abstract
Recent advances in computational pathology have led to the emergence of numerous foundation models. However, these approaches fail to replicate the diagnostic process of pathologists, as they either simply rely on general-purpose encoders with multi-instance learning for classification or directly apply multimodal models to generate reports from images. A significant limitation is their inability to emulate the diagnostic logic employed by pathologists, who systematically examine slides at low magnification for overview before progressively zooming in on suspicious regions to formulate comprehensive diagnoses. To address this gap, we introduce CPathAgent, an innovative agent-based model that mimics pathologists’ reasoning processes by autonomously executing zoom-in/out and navigation operations across pathology images based on observed visual features. To achieve this, we develop a multi-stage training strategy unifying patch-level, region-level, and whole-slide capabilities within a single model, which is essential for mimicking pathologists, who require understanding and reasoning capabilities across all three scales. This approach generates substantially more detailed and interpretable diagnostic reports compared to existing methods, particularly for huge region understanding. Additionally, we construct an expert-validated PathMMU-HR², the first benchmark for huge region analysis, a critical intermediate scale between patches and whole slides, as diagnosticians typically examine several key regions rather than entire slides at once. Extensive experiments demonstrate that CPathAgent consistently outperforms existing approaches across three scales of benchmarks, validating the effectiveness of our agent-based diagnostic approach and highlighting a promising direction for the future development of computational pathology.
1 Introduction
Figure 1: Comparison of the "black box" traditional MIL approach, LMM approach, and our proposed CPathAgent for analyzing pathology images. CPathAgent interpretably mimics pathologists’ reasoning by performing operational actions (e.g., zooming in, moving the view) while describing analytical logic.
Pathology serves as the gold standard for diagnosing numerous diseases, particularly cancer. Advances in digital pathology have enabled the digitization and computational analysis of histopathological slides, creating promising opportunities for AI-assisted diagnostics. Recently, various pathology foundation models have emerged to meet the growing demand for automated analysis in this field.
Current foundation models in computational pathology primarily fall into two categories (as shown in the left part of 1): encoder-based models like DINO[1; 2] or CLIP [3] that extract patch-level representations from Whole Slide Images (WSIs) followed by Multi-Instance Learning (MIL) [4; 5; 6] aggregation to perform diagnosis, and large multimodal models (LMMs) that process both visual data and text instructions to generate pathology reports or other outputs [7; 8; 9; 10]. Despite their impressive benchmark performance, these approaches share a critical limitation: they fail to incorporate the intermediate reasoning processes that expert pathologists employ, limiting their interpretability and clinical applicability.
In clinical practice, pathologists employ a methodical approach to slide examination, starting with low-magnification scanning (typically 2-4×) to assess tissue architecture and overall morphology. They then strategically navigate to regions of interest showing atypical features or diagnostic importance, progressively increasing magnification (10×, 20×, and 40×) to evaluate fine-grained histopathological features. Throughout this process, pathologists continuously shift between different magnifications and fields of view, integrating spatial and morphological information across multiple scales to formulate a comprehensive diagnosis. Current computational models bypass this sophisticated visual reasoning process, resulting in: (1) insufficient interpretability, as they cannot demonstrate visual navigation and multi-scale reasoning; (2) vulnerability to hallucinations and shortcut learning, potentially fixating on spurious correlations rather than relevant histological patterns; and (3) poor alignment with clinical workflows, undermining their credibility in diagnostic settings.
Drawing inspiration from recent breakthroughs in agent-based systems [11; 12; 13], which have demonstrated remarkable capabilities across diverse domains including strategic gameplay [14], web navigation [15], complex task automation [16], and multimodal content generation [17], we introduce CPathAgent, a novel framework that emulates the adaptive visual reasoning of expert pathologists. Unlike existing computational pathology models that operate as black-box classifiers or descriptors, CPathAgent implements a dynamic, perception-driven approach that mirrors pathologists’ hierarchical examination strategy in clinical practice. CPathAgent performs strategic operations: zooming in to examine cellular details, zooming out to maintain contextual awareness, and navigating across regions of interest, mimicking how pathologists scan slides at low magnification to identify suspicious areas before zooming in to examine fine-grained histological patterns before making diagnosis. This represents a fundamental paradigm shift in computational pathology from static, outcome-focused models to interactive, reasoning-based systems with greater clinical utility.
Specifically, CPathAgent mimics pathologists’ diagnostic workflows in a multi-stage process, as shown in the right part of Figure 1: (1) Global screening and region selection: CPathAgent initially analyzes a low-resolution WSI view to identify regions of interest (16000×16000 region at 40×) exhibiting suspicious histological features for detailed examination. (2) Navigation planning: For each prioritized region, CPathAgent plans a navigation path defining a sequence of view coordinates at varying magnifications, determining when to pan, zoom in, or zoom out based on the observed features within the huge region. (3) Multi-scale multi-view sequence reasoning: CPathAgent executes the planned navigation path by analyzing specific fields of view at different magnifications, then synthesizes observations across these views to perform step-by-step diagnostic reasoning.
Our work makes four significant contributions to advance agent-based computational pathology:
1) We introduce CPathAgent, which is a agent-based framework that emulates a pathologist’s diagnostic workflow through dynamic navigation planing and multi-magnification multi-view analysis, enabling interpretable, reasoning-based assessments.
2) We develop PathMMU-HR², the first visual question answering benchmark specifically designed for high-resolution huge pathology image regions (16000×16000 pixels), rigorously validated by three professional pathologists to ensure clinical relevance and diagnostic accuracy.
3) We introduce a multi-stage training strategy: An approach that equips CPathAgent with cognitive and reasoning capabilities across different scales, from patch-level details to whole-slide contexts, within a unified diagnostic framework.
4) We conduct extensive evaluations across benchmarks at multiple scales, ranging from patches to large regions to WSIs, highlighting CPathAgent’s strong performance and potential clinical utility.
2 Related Work
2.1 Pathology Language-Vision Models
Recent years have witnessed substantial progress in developing foundation models specifically tailored for computational pathology. These models broadly fall into two categories: encoder-based models and large multimodal models (LMMs).
Encoder-based models primarily aim to extract patch-level representations from whole-slide images (WSIs). Early efforts in this area relied on ImageNet-pretrained convolutional neural networks (CNNs) [18; 19] as feature extractors, combined with multi-instance learning (MIL) techniques [4; 5; 6] to aggregate these features for slide-level prediction. More recently, the field has shifted toward pathology-specific foundation models that adopt self-supervised learning paradigms. Architectures such as DINO [1; 2] and CLIP [3], when trained on large-scale pathology datasets, have demonstrated superior feature extraction capabilities [20; 8; 21; 22; 23]. These improved representations, integrated via MIL or incorporated into pretrained WSI transformers, have significantly advanced the state of the art in WSI-level classification. Large multimodal models (LMMs) for pathology have emerged in parallel, enabling the processing of pathology images alongside text-based instructions to generate textual outputs. These models integrate visual and linguistic information through various designs. Patch-level MLLMs [7; 24; 8; 25; 26] adapt general vision-language frameworks like LLaVA [27] to interpret pathology image patches, while WSI-level models [9; 28; 29] aim to process entire whole-slide images for generating reports or answering diagnostic questions. Recent unified frameworks [10] attempt to handle both patches and WSIs within a single model architecture.
Despite these advances, current models typically process pathology inputs in a static and fixed manner, which contrasts with the dynamic, interactive diagnostic workflow of expert pathologists, who routinely navigate across regions of interest and examine tissues at multiple magnification levels, integrating contextual information over time to formulate a diagnosis.
2.2 Agent-based Models
Agent-based models represent an emerging paradigm in AI research, characterized by systems that can perceive their environment, make decisions, and take actions to achieve specific goals [30]. Recent advances in large language models (LLMs) have catalyzed significant progress in this domain, enabling more sophisticated reasoning and planning capabilities.Notable examples include WebGPT [15] and WebShop [31], which navigate web environments to complete information-seeking and shopping tasks; ReAct [32], which interleaves reasoning and action steps for improved task performance; AutoGPT [33] and AutoGen [34], which decompose complex tasks into manageable subtasks with minimal human supervision; and task-specific agents for code development [17], creative content generation [35], scientific research [36], and clinical decision-making and discovery that simulate physician workflows in hospital environments [37; 38; 39; 40]. These agent models typically incorporate several key components [11]: perception modules to process environmental inputs, reasoning modules to interpret observations and plan actions, and execution modules to implement decided actions. This architecture enables agents to adapt to dynamic environments and accomplish complex, multi-step tasks through iterative observation-reasoning-action cycles.
Although agent-based approaches have demonstrated success across various domains, their application to computational pathology remains largely unexplored. PathFinder [41] represents a notable early attempt at introducing agent-based systems for WSI analysis. The system segments WSIs into fixed 512×512 patches and employs a U-Net-based architecture to generate importance maps. To be specific, PathFinder feeds WSI thumbnails to generate initial importance maps, selects high-probability regions for description generation, and regenerate subsequent importance maps based on previous descriptions and thumbnails. While PathFinder introduces agent-based concepts to WSI analysis, it primarily operates through region selection at a fixed magnification level within predefined 512×512 patches. Pathologists, however, employ a more flexible diagnostic workflow with dynamic multi-scale navigation—beginning with low-magnification overviews and continuously adjusting zoom levels while moving across tissue regions based on morphological observations to integrate information across different scales.
To address this gap, our CPathAgent takes a step further by incorporating multi-scale navigation capabilities that mimic pathologist behavior. It autonomously decides where to examine, when to zoom in or out, and how to navigate across WSIs, closely mimicking the iterative diagnostic workflow of expert pathologists. Notably, CPathAgent provides real-time verbalization of its diagnostic reasoning, effectively simulating the internal thought processes that pathologists experience during slide examination (specific examples can be found in Appendix Figure 16 to 20). This approach enables enhanced interpretability through traceable reasoning paths and improved diagnostic performance.
2.3 Pathology Benchmarks and Datasets
Benchmark datasets play a crucial role in advancing computational pathology by providing standardized evaluation frameworks. Traditional whole-slide image (WSI) benchmarks like Camelyon16 [42] and TCGA [43] focused primarily on slide-level classification tasks. More recently, datasets like WSICaption[44] and WSI-VQA[45] have emerged, targeting whole-slide image description generation and question-answering tasks across entire slide images, respectively.
In parallel, several patch-level benchmarks have emerged for evaluating multimodal models in pathology. PathVQA [46] introduced the first visual question answering (VQA) tasks for pathology images, drawing on textbook image -caption pairs. QuiltVQA expanded this approach by extracting image-caption pairs from pathology lectures on YouTube, resulting in higher-quality QA pairs due to the educational nature of these materials. PathMMU [47] marks a more recent advance, offering a comprehensive benchmark across diverse pathology sources. Its key strength is the expert validation process, resulting in a professionally curated dataset that better reflects clinical practice.
Despite recent progress, a critical gap remains in existing benchmarks: they focus primarily on either small, localized patch-level tasks or broad whole-slide analysis, overlooking an essential intermediate scale, the huge region, which plays a central role in real-world diagnostic reasoning. In routine practice, pathologists typically begin by scanning the whole slide images to identify suspicious areas, and then concentrate their diagnostic attention on large, yet manageable regions containing the relevant pathology. It is often within these huge regions that key diagnostic decisions are made, as they provide sufficient context for interpretation.
To address this limitation, we introduce PathMMU-HR², a benchmark specifically designed for huge region analysis in pathology. By targeting this intermediate scale—situated between isolated patches and WSIs, PathMMU-HR² better aligns with the way pathologists actually examine and interpret tissue. It provides a more controlled, context-rich, and diagnostically meaningful setting for evaluating multimodal models, especially those incorporating agent-based reasoning across different scales. To ensure clinical relevance, we also engaged three board-certified pathologists to validate the dataset, making PathMMU-HR² an expert benchmark for intermediate-scale pathology evaluation.
3 Methods
Figure 2: Illustration of CPathAgent framework, which mimics the diagnostic workflow of pathologists via global screening, navigation planning, and multi-scale reasoning across sequential views.
In this section, we provide a comprehensive description of our proposed approach. We detail the construction of CPathAgent-Instruct, our instruction-tuning dataset used to train CPathAgent, the development of the PathMMU-HR2 benchmark dataset, and the specific training methodology implemented for the CPathAgent model.
3.1 Overview of CPathAgent Reasoning Workflow
To simulate expert-level diagnostic reasoning, we design CPathAgent as an agent-based system capable of dynamic region selection, strategic navigation planning, and multi-scale reasoning. As illustrated in Figure 2, the entire diagnostic process is decomposed into three key stages:
1) Global Screening: CPathAgent first scans the low-resolution thumbnail of a whole-slide image to identify and cluster suspicious regions. Uninformative areas are filtered out, while large diagnostically relevant regions (16000$\times$16000 pixels) are preserved for further analysis.
2) Navigation Planning: For each preserved huge region, the model plans a navigation path across spatial coordinates and magnification levels (e.g., 1$\times$, 3$\times$, 4.5$\times$). The goal is to plan sequentially examine areas that are most likely to yield diagnostic insights.
3) Multi-scale Multi-view Sequence Reasoning: Along the planned navigation path, the model observes each field of view, integrates visual features across scales and generates step-by-step reasoning descriptions similar to a pathologist’s thought process explaining the purpose of actions like zooming in and moving and finally synthesizes findings into a coherent diagnostic conclusion.
3.2 CPathAgent-Instruct Dataset Construction
To effectively train our agent-based model to simulate the complex diagnostic reasoning of expert pathologists, we construct the CPathAgent-Instruct dataset, which consists of several subsets targeting different critical reasoning capabilities across multiple scales.
Unlike approaches that rely solely on prompt engineering during inference, our focus is on generating high-quality instruction data to train open-source models capable of expert-level pathology reasoning. The data generation strictly follows the reasoning workflow illustrated in Figure 1 but leverages a additional information to enhance data quality, summarized as follows:
Region Selection Dataset: We start by downscaling each WSI by a factor of 32 to create navigable thumbnails. We define huge regions (16000×16000 pixels at 40× magnification) with 5% overlap between adjacent regions to preserve continuity across region boundaries. We prompt Gemini-2.5-Pro with carefully crafted prompts and WSI paired reports (detailed in Appendix D) to generate structured region annotations that include: (1) region groupings with descriptive names, (2) prioritized ID lists representing recommended viewing order, and (3) flags indicating necessity for high-magnification examination. This process produces thumbnail-region selection pairs that emulate a pathologist’s systematic WSI screening strategy.
Viewing Path Planning Dataset: To teach the model strategic examination sequences across multiple magnifications, we implement a three-step reference-guided approach: (1) We extract region-specific content from expert-written WSI reports by prompting Gemini-2.5-Pro with specific region and WSI report (upper left panel of Figure 3); (2) We first crop patches at three magnification levels (1× for comprehensive overview, 2× for four quadrant views, and 4× for sixteen high-resolution details), then send both these multi-scale patches and the previously extracted region-specific captions to Gemini-2.5-Pro, this guidance from expert descriptions enables the model to generate more accurate and diagnostically relevant patch descriptions (upper right panel of Figure 3); Having detailed descriptions at different scales guides the model to more effectively generate navigation planning; (3) We then annotate images with coordinate grids (0.1 relative position intervals) as reference, and use the previously generated patch descriptions to prompt Gemini-2.5-Pro to produce viewing paths. We specifically require generation of variable-scale paths that can navigate to any scale and position, rather than fixed-scale routes. These steps produce paired huge region-navigation planned path data that forms the foundation of our viewing path dataset.
Multi-Image Understanding and Reasoning Dataset: We extract image crops at specified magnification levels along the generated viewing paths (lower right panel of Figure 3). We pair these multi-scale image sets with WSI report fragments and prompt Gemini-2.5-Pro to generate step-by-step visual reasoning sequences. The resulting dataset contains structured reasoning chains that integrate observations across magnification levels, identify specific morphological features, and formulate diagnostic interpretations based on collective evidence, enabling CPathAgent to develop sophisticated multi-scale reasoning capabilities.
Figure 3: Overview of the generation process for CPathAgent’s navigation planning subset and multi-scale multi-view sequence reasoning instruction-tuning data.
VQA-oriented Dataset: To enable CPathAgent to generate question-oriented dynamic viewing paths, we develop a specialized dataset for huge region VQA. First, we generate diverse clinical questions based on previously generated multi-scale patch descriptions. Next, we adapt our viewing path planning approach by prompting Gemini-2.5-Pro with these questions alongside region descriptions to generate examination paths optimized for answering each clinical query. Finally, we synthesize step-by-step reasoning chains that navigate through relevant visual evidence to reach clinically sound answers. Unlike our previous datasets guided by comprehensive region descriptions, this dataset is driven by specific visual questions, training CPathAgent to adapt its examination strategy based on question context—selecting optimal magnification levels and fields of view to efficiently answer diagnostic queries.
Overall, we generated 278K instruction-tuning data points that collectively capture the dynamic, multi-scale reasoning processes that mimic expert pathologists’ diagnostic workflows. The resulting CPathAgent-Instruct dataset equips our model with advanced capabilities essential for computational pathology: strategic region selection, examination planning, cross-magnification feature integration, and targeted diagnostic reasoning and question answering. For more detailed dataset statistical information, please kindly refer to Appendix A.
3.3 PathMMU-HR2 Dataset Construction
To validate CPathAgent’s capabilities in huge region analysis, we developed the PathMMU-HR2 benchmark dataset. We ensured diverse pathological representation by sampling regions across various tissue types (e.g., TCGA-BRCA, TCGA-LUAD, etc.) from The Cancer Genome Atlas (TCGA), and selecting ultra-regions from different histological groups (as detailed in section 3.2) to capture a wide range of pathologically significant structures. Using an approach similar to CPathAgent-Instruct development, we generated VQA pairs that specifically require integration of features across multiple magnification levels. Three board-certified pathologists rigorously evaluated all generated VQA pairs. Their assessment focused on the clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration to answer questions, and alignment with standard pathology practice. VQA pairs that failed to meet any of the established quality criteria were eliminated from the dataset. The finalized PathMMU-HR2 benchmark contains 1,688 expert-validated VQA instances specifically designed to evaluate multi-scale reasoning capabilities in huge regions.
Figure 4: Overview of architecture and three-stage training pipeline of CPathAgent.
3.4 CPathAgent Model Architecture and Training
CPathAgent employs a multimodal architecture built upon the LLaVA-OneVision [48] framework, with significant adaptations tailored for computational pathology tasks. At its core, CPathAgent integrates two powerful components: Qwen3-14B [49], a powerful LLM that provides strong reasoning and generation capabilities, and CPath-CLIP, a powerful pathology-specific CLIP model trained on over 700,000 high-quality pathology image-caption pairs. These components are connected through a two-layer MLP projection layer.
We implement a multi-stage training strategy, as shown in Figure 4 which is a progressive three-stage approach that systematically builds CPathAgent’s capabilities from foundational multimodal alignment to advanced agent-based diagnostic reasoning:
Stage 1 focuses exclusively on aligning the feature spaces between CPath-CLIP and the Qwen3-14B LLM. Only the two-layer MLP projection layer connecting the vision encoder to the language model is trained, while keeping both base models frozen. This training utilizes the CPath-PatchCaption dataset [10], enabling effective pre-alignment of visual and language representations without disrupting the pre-trained capabilities of either component.
Stage 2 builds upon the aligned feature spaces from Stage 1, where we train CPathAgent on comprehensive pathology tasks using the CPath-PathInstruct dataset [10]. During this stage, all model parameters are unfrozen for fine-tuning, allowing CPathAgent to master diverse patch-level tasks such as visual question answering, image classification, captioning, and diagnostic reasoning. This stage establishes the foundation for CPathAgent’s higher-level diagnostic capabilities.
Stage 3, the most crucial phase, focuses on transforming CPathAgent from a multimodal model into an agent capable of emulating pathologists’ diagnostic workflows. We combined our specialized CPathAgent-Instruct dataset with 20% of CPath-Instruct for training. This integration equips CPathAgent with agent capabilities that mimic pathologists’ whole-slide image navigation and large region reasoning processes, while preserving the strong patch-level understanding developed in previous stages. As in Stage 2, all model parameters remain unfrozen during this stage.
Our multi-stage training strategy produces a versatile model that excels at both detailed patch-level analysis and strategic large-region and whole-slide image navigation and examination, making CPathAgent a powerful tool that can adapt to various clinical scenarios requiring different levels of pathological analysis.
4 Experiments
We evaluate CPathAgent across multiple tasks: patch understanding, huge region analysis, and WSI classification to assess its capabilities and benchmark performance against existing approaches.
4.1 Patch Understanding Evaluation
Since patch-level understanding serves as the foundation for reasoning over larger regions and for planning across WSIs, we first evaluate CPathAgent’s ability to interpret standard-resolution pathology image patches. We conduct this evaluation on PathMMU, currently the largest and most diverse expert-verified pathology dataset. We compare CPathAgent with the most advanced general-purpose LMMs, including InstructBLIP-FLAN-T5 XXL[50], LLaVA-1.5-13B[51], LLaVA-OneVision[48] series, Qwen2.5-VL[52] series, GPT-4V[53], GPT-4.1[54] series and Gemini-2.5[55] series, along with domain-specific pathology LMMs such as LLaVA-Med[56], Quilt-LLaVA[8], PathGen-LLaVA[26] and CPath-Omni[10].
Results: CPathAgent significantly outperforms both general-purpose and pathology-specific LMMs across all subsets. As shown in Table 1, it achieves the highest performance across all subsets, with 80.5% on the tiny test set and 78.6% on the full test set. This shows a clear margin over the strongest general-purpose model (Gemini-2.5-Pro, 68.7% Tiny / 67.5% All) and surpasses the SOTA domain-specific model (CPath-Omni, 72.4% Tiny / 72.2% All). While part of CPathAgent’s patch-level performance can be attributed to the training data inherited from CPath-Omni, CPathAgent benefits further from the third-stage training with multi-scale and reasoning-focused data. This helps it develop stronger reasoning ability and builds a solid foundation for simulating expert-like behaviors such as zooming, navigation, and diagnostic thinking across WSIs.
CPathAgent even outperforms the expert-annotated baselines in several subsets, suggesting that its general patch-level understanding is already quite promising. This may reflect limitations in human annotations, as individual experts often specialize in narrow domains and may struggle with unfamiliar cases in a diverse dataset like PathMMU. While CPathAgent benefits from broad training and shows potential as a general-purpose pathology model, patch-level tasks remain relatively simple—real clinical diagnosis often demands more complex reasoning across larger spatial contexts.
Table 1: Overall results of models on the PathMMU test set. The best-performing LMM in each subset for general and pathology LMMs is in-bold, and the second-best performing LMM is underlined.
Test Overall PubMed SocialPath EduContent Atlas PathCLS
Tiny ALL Tiny ALL Tiny All Tiny All Tiny ALL Tiny ALL
(1156) (9677) (281) (3068) (235) (1855) (255) (1938) (208) (1007) (177) (1809)
Expert performance 71.8 - 72.9 - 71.5 - 69.0 - 68.3 - 78.9 -
General Large Multimodal Models
InstructBLIP-FLAN-T5-XXL 34.3 33.9 39.1 37.2 33.6 34.3 34.5 36.0 38.5 39.3 22.6 22.7
LLaVA-1.5-13B 38.8 37.6 44.5 41.0 40.4 40.4 34.1 39.4 47.1 44.3 24.9 23.5
LLaVA-OneVision-Qwen2-7B-OV 36.9 34.4 37.7 36.4 35.7 38.4 47.1 38.3 38.9 38.4 20.3 20.4
LLaVA-OneVision-Qwen2-72B-OV 51.0 46.4 60.5 51.0 59.1 52.9 54.9 49.8 43.8 49.2 27.7 26.5
Qwen2.5-VL-7B-Instruct 39.7 37.5 40.2 40.2 41.7 39.0 47.1 42.0 40.4 38.9 24.9 25.6
Qwen2.5-VL-72B-Instruct 56.2 51.2 63.3 56.1 62.4 55.5 65.5 58.3 54.3 55.2 26.0 28.7
GPT-4V-1106 53.9 49.8 59.4 53.5 58.7 53.9 60.4 53.6 48.1 52.8 36.2 33.8
GPT-4.1-mini-2025-04-14 60.7 59.9 66.9 63.4 62.1 61.8 65.1 61.1 60.1 62.8 43.5 48.8
GPT-4.1-2025-04-14 67.7 64.4 73.0 66.9 70.6 65.5 69.4 64.9 64.9 66.8 56.5 57.2
Gemini-2.5-Flash-Preview-04-17 68.0 65.2 74.9 69.2 71.1 64.9 68.1 65.3 67.5 69.3 53.7 56.3
Gemini-2.5-Pro-Preview-03-25 68.7 67.5 73.8 71.5 72.1 68.1 70.7 68.6 68.3 70.6 53.7 57.5
Pathology-specific Large Multimodal Models
LLaVA-Med 25.3 26.2 28.5 27.7 28.9 27.3 22.7 27.2 22.6 30.7 22.6 20.3
Quilt-LLaVA 45.6 41.5 47.3 42.6 46.4 46.6 51.8 45.3 46.2 42.7 32.2 29.2
PathGen-LLaVA 60.1 58.4 60.1 60.1 60.9 58.8 60.8 60.7 63.5 64.9 54.2 48.9
CPath-Omni 72.4 72.2 74.0 69.9 76.6 71.8 69.8 70.6 65.9 70.6 75.7 79.0
CPathAgent (Ours) 80.5 78.6 80.8 78.4 83.4 77.9 83.5 79.6 75.4 77.6 78.0 79.2
4.2 Huge Region Understanding Evaluation
We evaluate CPathAgent’s ability to autonomously navigate and reason over huge, high-resolution pathology regions using our PathMMU-HR² benchmark, which includes 1,688 expert-validated VQA pairs requiring multi-scale analysis. This evaluation is designed to assess whether models can mimic expert-like behaviors—such as zooming in and out, adjusting magnification, and selectively focusing on diagnostically relevant areas. Similar to Section 4.1, we involve the same set of models, but one difference is that we apply the prompt used by CPathAgent, enabling stronger agent-capable models like Gemini-2.5-Pro and GPT-4.1-mini to directly mimic the pathologist’s zoom-in/zoom-out movement reasoning process to assess whether this can further enhance their capabilities. In addition, we also invited a pathology expert to evaluate both the diagnostic description generation tasks and VQA tasks for 40 regions, including whether their navigation path planning and multi-scale multi-view reasoning aligns with clinical logic. We ran the model 8 times to calculate the pass@k.
Results: CPathAgent demonstrates superior performance on huge, high-resolution pathology regions across diverse cancer types. As shown in Table 2, CPathAgent achieves an impressive 88.6% accuracy on the PathMMU-HR² benchmark, substantially outperforming all other models. Despite having a more limited single-view capability (1008×1008 resolution) compared to larger models like Gemini-2.5-Pro (3072×3072 resolution), CPathAgent effectively overcomes this limitation by mimicking pathologists’ diagnostic workflow, strategically navigating, zooming in on relevant regions, and integrating observations across multiple views to form comprehensive diagnostic reasoning. We provide examples of CPathAgent’s planning and navigation process in Appendix D.
Agent-based approaches consistently enhance model performance for pathology analysis. As shown in Table 2, the results clearly demonstrate that incorporating agent-like navigation capabilities improves performance across models. Overall, models equipped with CPathAgent’s prompting strategy show an average performance improvement of 3.2 percentage points, with Gemini-2.5-Pro improving from 73.2% to 76.4% and GPT-4.1-mini from 60.1% to 62.3% on the full benchmark. The enhancement is particularly pronounced in complex cancer types like KICH, where Gemini-2.5-Pro shows a dramatic improvement from 56.5% to 69.3% (+12.8%). This consistent pattern across models and cancer types confirms that the ability to dynamically explore and reason over pathology images in a manner similar to expert pathologists is crucial for accurate diagnosis, especially for cases requiring integration of findings across multiple scales and regions. These promising results suggest that empowering vision models with expert-like navigation strategies could be a key approach for advancing AI performance in complex medical imaging tasks beyond pathology.
Pathologist evaluation confirms that the navigation paths and diagnostic reasoning generated for both tasks meet clinical expectations. As shown in 5, VQA-based approaches achieve higher satisfaction rates than general description tasks. We attribute this to VQA’s focused attention on target features, enabling better identification of key elements, while general descriptions require additional judgment regarding which content should be highlighted first. Furthermore, as shown in Figure 5, the pass@k metric demonstrates an upward trend as the number of model runs increases, indicating the model’s potential to achieve improved results through multiple generations.
Table 2: Performance comparison of huge region classification across multiple cancer types using two approaches: General LMMs, and Agent-based Methods. The best-performing model for each cancer type is in-bold, and the second-best performing model is underlined.
BRCA LUAD LUSC KIRP KIRC KICH ESCA THCA BLCA TGCT Overall
(368) (192) (230) (153) (151) (62) (70) (249) (141) (52) (1688)
General LMM Approach
LLaVA-OneVision-Qwen2-7B-OV 42.9 43.2 33.5 28.1 37.1 43.5 40.0 41.8 39.7 30.8 38.8
LLaVA-OneVision-Qwen2-72B-OV 40.8 52.1 40.0 39.2 36.4 33.9 35.7 45.8 57.4 36.5 43.0
Qwen2.5-VL-7B-Instruct 47.3 57.3 35.2 46.4 47.7 25.8 51.4 49.8 56.0 38.5 46.9
Qwen2.5-VL-72B-Instruct 48.4 61.5 44.3 52.9 56.3 27.4 52.9 48.6 62.4 51.9 51.2
GPT-4.1-mini-2025-04-14 56.0 64.6 60.4 62.1 61.6 35.5 70.0 59.4 68.1 59.6 60.1
GPT-4.1-2025-04-14 57.9 63.5 62.2 67.3 69.5 54.8 65.7 66.3 67.4 71.2 63.7
Gemini-2.5-Flash-Preview-04-17 66.6 74.0 76.1 65.4 76.8 40.3 70.0 72.7 76.6 63.5 70.4
Gemini-2.5-Pro-Preview-03-25 64.4 70.8 78.3 74.5 80.8 56.5 77.1 76.7 80.1 75.0 73.2
Quilt-LLaVA 29.3 45.8 44.8 38.6 38.4 30.6 35.7 39.8 47.5 40.4 38.8
PathGen-LLaVA 62.0 77.6 67.4 58.2 74.2 48.4 67.1 84.4 48.1 67.2 67.2
Agent-based Approach
GPT-4.1-mini-2025-04-14 56.0 69.3 63.9 62.7 64.9 41.9 71.4 59.8 68.1 73.1 62.3
Gemini-2.5-Pro-Preview-03-25 68.8 68.8 80.0 77.8 83.4 69.3 82.8 78.7 85.1 76.9 76.4
CPathAgent (Ours) 87.0 88.5 87.8 87.9 92.9 78.9 90.7 89.0 90.7 93.0 88.6
Figure 5: Success rates (pass@k) across 8 generations, evaluated by pathologists across 40 regions. Results show the percentage of pathologist satisfaction with navigation paths and reasoning for General Description Generation (left) and VQA (right) as the number of generations increases.
4.3 WSI Classification
We evaluate CPathAgent’s performance on Whole Slide Image (WSI) classification across six diverse datasets spanning multiple cancer types and diagnostic tasks: TCGA-BRCA (breast cancer subtyping), TCGA-NSCLC (lung cancer subtyping), TCGA-RCC (renal cell carcinoma subtyping), TCGA-ESCA (esophageal cancer subtyping), TCGA-BLCA (bladder cancer subtyping), and TCGA-THCA (thyroid cancer subtyping). We benchmark CPathAgent against traditional MIL approaches, including ABMIL and DSMIL, as well as both general-purpose and pathology-specific multimodal foundation models.
Since CPathAgent generates detailed diagnostic descriptions rather than simple classification labels, we implement a two-stage evaluation process to fairly compare with existing benchmarks. First, CPathAgent employs our agent-based approach—scanning at low magnification to identify suspicious regions, then strategically analyzing specific areas to generate comprehensive diagnoses for all significant regions within each WSI. These detailed descriptions capture the complex pathological findings but must be translated to match the predefined classification schemes of standard benchmarks. Therefore, we provide these comprehensive descriptions to Gemini-2.5-Pro for classification into the specific labels required by each dataset. The more accurate and detailed CPathAgent’s descriptions, the more precise the resulting WSI classifications.
Results: CPathAgent achieves competitive performance on whole slide image classification tasks despite using a fundamentally different approach. As shown in Table 3, CPathAgent achieves an average accuracy of 82.8% across six different cancer classification tasks, outperforming traditional MIL approaches (ABMIL: 79.9%, DSMIL: 76.8%) and substantially surpassing general-purpose LMMs like Gemini-2.5-Pro (72.1%). CPathAgent demonstrates particularly strong performance on TCGA-BRCA (88.5%) and TCGA-ESCA (97.1%), where it matches or exceeds specialized MIL methods. While the overall margin over MIL approaches is modest (+2.9% vs. ABMIL), this represents a significant paradigm shift—CPathAgent achieves these results through interpretable, expert-like diagnostic reasoning rather than through opaque feature aggregation.
The agent-based approach demonstrates substantial headroom for further improvement. The performance gap between CPathAgent (82.8%) and the theoretical upper bound (91.7%) suggests significant potential for further optimization. This upper bound represents the performance achieved by directly using our generated expert-quality data for whole slide image classification, highlighting both the quality of our training data and the considerable margin still available for improvement. Despite using high-quality training data, CPathAgent still shows a nearly 9% gap to the overall upper bound, particularly in challenging datasets like TCGA-BLCA and TCGA-THCA. This suggests that enhancing agent navigation strategies and reasoning capabilities could yield substantial gains.
Table 3: Overall results of WSI classification tasks. The best-performing model in each subset for general and pathology domain is in-bold, and the second-best performing model is underlined. Balanced accuracy (%) is employed.
TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
Multi-instance Learning Approach (Traditional)
ABMIL 80.5 92.8 96.4 91.2 50.1 68.4 79.9
DSMIL 84.7 87.2 88.9 89.9 52.2 57.8 76.8
General Large Multimodal Models
GPT-4.1-mini-2025-04-14 52.8 65.2 52.8 85.3 57.0 54.3 59.6
GPT-4.1-2025-04-14 61.0 65.8 53.8 85.3 61.3 58.9 64.3
Gemini-2.5-Flash-Preview-04-17 56.7 84.5 57.9 93.8 61.8 55.4 68.4
Gemini-2.5-Pro-Preview-03-25 72.4 89.2 69.2 97.1 59.8 44.9 72.1
Agent-based Approach
CPathAgent (Ours) 88.5 90.8 94.6 97.1 62.7 63.2 82.8
[…]
5 Conclusion
In this study, we introduced CPathAgent, an agent-based approach for computational pathology that emulates pathologists’ diagnostic reasoning through strategic examination of high-resolution pathology images. By meticulously replicating critical diagnostic processes—such as zooming in to inspect cellular details, zooming out to assess tissue architecture, and systematically shifting focus between regions of interest, CPathAgent achieves both enhanced interpretability and state-of-the-art performance across diverse pathology tasks. Our comprehensive evaluation across three levels of image scale (patch-level understanding, large regional assessment, and whole slide image analysis) demonstrates that this pathology-mimicking approach, which closely mirrors the trained eye of experienced pathologists navigating tissue specimens, effectively bridges the crucial gap between human diagnostic expertise and computational analysis.
Supplementary Material
[…]
Appendix A Additional Details of Proposed Dataset
In this section, we present detailed information about our datasets: CPathAgent-Instruct and PathMMU-HR². We provide comprehensive statistical analysis of both datasets, along with further details regarding their construction process.
A.1 Details of CPathAgent-Instruct
Figure 6: Overview of the VQA-oriented dataset generation process: from multi-scale description to question generation, question filtering, question-oriented navigation path generation, and VQA-related visual reasoning.
Image Sources: CPathAgent derives its WSI report data from HistGen [57], while the corresponding whole slide image data is sourced from The Cancer Genome Atlas (TCGA). TCGA provides a comprehensive collection of pathology whole-slide images contributed by various participating institutions, encompassing a diverse range of distinct tissue and cancer types. This extensive repository ensures diverse coverage of pathological conditions and features essential for dataset and model development.
Additional Details of CPathAgent-Instruct: Our dataset utilizes 5,254 WSIs, which is 80% of the total HisGen data for constructing CPathAgent-Instruct. The remaining WSIs are reserved for model testing and for constructing the PathMMU-HR² dataset.
We provide the following additional dataset construction details that could not be included in the main text due to space constraints:
1) Region Selection Subset Enhancement: For the region selection subset within CPathAgent-Instruct, we face a challenge as this capability was not developed during earlier training stages (as detailed in Section C), and 5,254 WSIs are insufficient for acquiring this skill. Considering that region selection is manageable for advanced large models like Gemini-2.5-Pro and doesn’t heavily depend on WSI reports, we expand our data by utilizing all 24,429 TCGA overview images (excluding test data). The WSI overview images are 32× downsampled versions of the original WSIs as mentioned in the main paper, providing a overview the entire slide at lower resolution. To be specific:
  • For WSIs with associated pathology reports, we prompt Gemini-2.5-Pro with both the WSI overview image and the corresponding report. The prompt specifically guides the model in identifying and selecting suspicious regions, particularly those exhibiting pathological features explicitly mentioned in the report.
  • For WSIs without accompanying reports, we provide only the WSI overview image, prompting the Gemini-2.5-Pro to identify regions showing abnormal tissue characteristics or potential pathological features by leveraging Gemini-2.5-Pro’s powerful visual recognition capabilities.
To increase the diversity of our data, we use a generation temperature of 0.8 and create two separate region selection results for each WSI. This strategy produces slightly different region selections from the same WSI overview, effectively doubling our dataset size. At the same time, it introduces natural variations in how regions are selected, which reflects the inherent subjectivity in how different pathologists might evaluate the same slide. The specific prompt templates we use can be found in Section D.5.
2) Navigation Path Diversity Augmentation: Similarly, for the navigation path planning subset, we recognize that pathologists can follow multiple valid paths to reach accurate diagnoses when examining a large region. To capture this diversity, we set the generation temperature to 0.8 and prompt Gemini-2.5-Pro 1-3 times per WSI using the approach described in Section 3.2. This generates diverse navigation paths for the same region. Correspondingly, the multi-scale multi-image reasoning also changes based on the different navigation paths, creating additional variations in our dataset. As a result, we substantially expand our dataset while improving the model’s ability to generate diverse, valid outputs.
3) VQA-oriented Subset Construction and Quality Control: We systematically construct the VQA subset as shown in Figure 6. First, we use 21 multi-scale descriptions from sub-patches at different magnifications (1x, 2x, 4x), which are generated through WSI-report guided prompting (detailed in Section 3.2). We then input these descriptions along with their corresponding region images to Gemini-2.5-Pro to create the initial VQA pairs.
However, we recognize that models often exploit textual shortcuts when answering VQA questions rather than performing genuine image analysis. Inspired by PathMMU [47], we present text-only questions to both Gemini-2.5-Pro [55] and DeepSeekV3 [58], then eliminate any questions both models answer correctly without visual input. This ensures our dataset exclusively targets vision-dependent reasoning tasks that cannot be solved through textual cues alone.
Figure 7: Illustration of the proportional distribution of each subset within the CPathAgent-Instruct dataset.
Figure 8: Illustration of the distribution of regions across tissue types.
Figure 9: Illustration of the distribution of regions across tissue types.
For questions that pass this validation, we generate navigation plans by providing Gemini-2.5-Pro with the question, multi-scale image descriptions, region descriptions, and full region images. Similar to our region description reasoning generation, we combine the validated questions with cropped patches from the planned navigation path and region descriptions. This prompts Gemini-2.5-Pro to generate comprehensive reasoning that systematically addresses each question—progressing logically from overall tissue architecture to specific diagnostic features, comparing and eliminating answer options through continuous reasoning to arrive at the correct conclusion.
Through this question-guided navigation planning and visual reasoning approach, we create a dataset that enables CPathAgent to adapt to diverse diagnostic queries, which allows the model to focus on relevant regions and features based on the specific question.
Proportion of CPathAgent-Instruct subsets Figure 7 illustrates the distribution of different subsets within the CPathAgent-Instruct dataset. Among them, "Viewing Path Planning" and "Multi-Image Understanding and Reasoning" each constitute 32.8% of the total dataset. These equal proportions exist because each planning instance directly corresponds to a reasoning instance, creating a one-to-one relationship between these components. The VQA-oriented subset, though smaller, is sufficient for model adaptation as it shares structural similarities with the main components.
Statistics of Source Regions: As shown in Figure 8, CPathAgent dataset encompasses 26 distinct tissue types with a total of 78,658 individual regions. Breast tissue represents the highest proportion at 13.4% (10,521 samples), while lymphoid tissue appears with 270 samples.
Statistics of Overview WSIs within region selection subset: As shown in Figure 9, the CPathAgent dataset consists of 26 distinct tissue types with a total of 24,429 WSI overviews. Lung tissue represents the highest proportion at 12.9% (3,150 WSIs), followed by Breast with 3,110 WSIs and Kidney with 2,672 WSIs.
Figure 10: Visualization of a word cloud derived from general description-based and VQA-oriented reasoning.
Figure 11: Visualization of a word cloud derived from questions in VQA-oriented subsets in CPathAgent-instruct and PathMMU-HR².
Word Frequency in Multi-scale Multi-image Reasoning Subsets: We analyze word frequency distributions in our General description-based (left, Figure 10) and VQA-oriented (right, Figure 10) reasoning approaches. The visualization reveals distinctive reasoning patterns in pathology diagnosis. General description reasoning emphasizes procedural terms like "zoom," "move," and "examine"—simulating a pathologist’s physical navigation through tissue samples as they methodically adjust magnification, reposition focus areas, and explore specimens. Meanwhile, VQA-oriented reasoning prioritizes diagnostic decision terms like "look," "see," and "feature," reflecting hypothesis testing and evidence evaluation in clinical reasoning. Both approaches incorporate critical pathological terminology ("stroma," "nuclear," "infiltrate") essential for accurate diagnosis, demonstrating that these reasoning strategies center on key diagnostic elements fundamental to clinical pathology assessment and disease classification.
Word Frequency of Questions in VQA-oriented Subset: We analyze the question terminology in our VQA-oriented subset, as shown in the left part of Figure 11. The visualization reveals "diagnosis" as the central focus, surrounded by practical terms like "describe," "integrate," and "feature." Magnification terms ("40x," "10x") indicate how pathologists navigate between different viewing scales, while specific organs ("thyroid," "breast," "lung") appear frequently in questions. Action words like "evaluate" and "observe," along with uncertainty terms like "likely" and "consider," demonstrate the methodical reasoning process. This word pattern demonstrates that our constructed diagnostic questions demand multiscale examination with systematic feature assessment.
A.2 Details of PathMMU-HR²
The construction of PathMMU-HR² is similar to the construction of the VQA-oriented subset in CPathAgen-Instruct, as shown in Figure 6. However, they differ in that after filtering VQA pairs that can be solved through text-only shortcuts, we engage three pathologists with over 10 years of experience for human validation. The evaluation criteria include: clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration for answering, and alignment with standard pathology practice. Questions failing any criterion are deemed invalid. Under pathologists’ evaluation, we identify 1,688 valid questions from the initially annotated 2,200 questions, with the annotation platform interface shown in Figure 12.
We also visualize the question part of PathMMU-HR² with a word cloud in Figure 11. As shown, the most prominent terms—such as "diagnose," "integrate," "accurately," and "across"—highlight that many queries require multi-scale analysis and complex reasoning. The presence of micro-level descriptors like "nuclear," "cellular," and "morphology," alongside macro-level terms such as "tumor," "lesion," and "pattern" underscores the necessity of integrating observations across different magnification levels (indicated by terms like "40x," "10x," and "2x"). This multi-scale integration is essential for correctly answering questions that demand the model to synthesize information from subcellular details to tissue architecture, evaluating both cytological features and architectural patterns simultaneously to arrive at accurate answers.
Figure 12: Example of the PathMMU-HR2 annotation interface.
Appendix B Additional Experiments and Discussion
B.1 Comparison of Description Generation Performance Among Pathology-specific LMMs
We randomly sample 100 region samples from TCGA and compare them with two advanced pathology-specific trained LMMs, PathGen-LLaVA [26] and Quilt-LLaVA [59], to generate descriptions for these regions. To assess model performance, we utilize Gemini-2.5-Pro and GPT-4.1 as evaluators, with outputs categorized as wins, losses, or ties when performance is comparable. To ensure clinical relevance of the evaluation, we also engage a professional pathologist with over 10 years of clinical experience to provide expert assessment.
Results: CPathAgent significantly outperforms both PathGen-LLaVA and Quilt-LLaVA across all evaluation metrics. As shown in Figure 13, most notably, CPathAgent achieves a perfect 100% win rate against Quilt-LLaVA across all three evaluators. Against PathGen-LLaVA, CPathAgent maintains consistently strong performance with over 95% win rates across all evaluators.
The evaluation consistency between human and LMM evaluators is remarkably high. However, we observe subtle differences in evaluation patterns: compared to the general-purpose LMMs, the pathologist evaluator awards PathGen-LLaVA slightly more ties and one additional win. We hypothesize this difference stems from the distinct evaluation priorities between expert pathologists and LMMs. General-purpose LMMs, may favor CPathAgent’s comprehensive reasoning, while the pathologist evaluator places equal weight on both reasoning quality and diagnostic accuracy, resulting in a more reliable assessment that recognizes PathGen-LLaVA’s clinical precision in some cases.
Figure 13: Comparison of CPathAgent with different pathology-specific LMMs for description/diagnosis generation on 100 randomly sampled cases, evaluated by Gemini-2.5-Pro, GPT-4.1, and expert pathologists.
B.2 Comparison of WSI Classification Performance Among Pathology-specific LMMs
We compare CPathAgent with existing pathology-specific LMMs on WSI classification tasks, following the same approach as used for general LMMs detailed in Section 4.3.
Results: Both PathGen-LLaVA and Quilt-LLaVA significantly underperform on WSI classification, achieving only 60.2% and 54.8% overall balanced accuracy, respectively, as shown in Table 4. This performance gap can be attributed to two key limitations of these models. First, PathGen-LLaVA and Quilt-LLaVA are specifically designed and trained for patch-level analysis using small tissue regions. Consequently, they lack the ability to capture high-level, contextual diagnostic information that is crucial for WSI analysis. Second, these models do not possess the capability to navigate and zoom within large-scale regions. When large regions are resized to fit their input requirements, critical diagnostic details become virtually indiscernible, severely compromising their performance on WSI classification tasks.
Table 4: Overall results of WSI classification tasks. The best-performing model in each subset for general and pathology domain is in-bold, and the model with the second-best performance is underlined. Balanced accuracy (%) is employed.
TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
General Large Multimodal Models
GPT-4.1-mini-2025-04-14 52.8 65.2 52.8 85.3 57.0 54.3 59.6
GPT-4.1-2025-04-14 61.0 65.8 53.8 85.3 61.3 58.9 64.3
Gemini-2.5-Flash-Preview-04-17 56.7 84.5 57.9 93.8 61.8 55.4 68.4
Gemini-2.5-Pro-Preview-03-25 72.4 89.2 69.2 97.1 59.8 44.9 72.1
Pathology-specific Large Multimodal Models
Quilt-LLaVA 55.3 54.0 53.3 58.6 61.9 45.8 54.8
PathGen-LLaVA 66.5 53.5 59.0 78.6 54.7 49.2 60.2
Agent-based Approach
Gemini-2.5-Pro-Preview-03-25 52.0 73.4 58.8 97.1 64.7 46.3 65.4
CPathAgent (Ours) 88.5 90.8 94.6 97.1 62.7 63.2 82.8
[…]
B.3 Ablation Study on Applying Gemini-2.5-Pro as an Agent for WSI Tasks
We also apply CPathAgent’s inference prompts to directly prompt Gemini-2.5-Pro, enabling it to perform agent-based reasoning and diagnosis similar to CPathAgent.
Results: Interestingly, while this approach yields modest improvements on TCGA-BLCA and TCGA-THCA datasets, the overall performance decreases by 6.7% (from 72.1% to 65.4%), as shown in the "Agent-based Approach" part of Table 4. This contrasts with the findings in Section 4.2 of the main paper, where applying agent-based approaches to advanced LMMs on large regions VQA demonstrated performance gains.
We attribute this discrepancy to the fundamental differences between task types. Since our WSI classification task involves first generating descriptions for all regions and then performing classification based on these descriptions, unlike VQA tasks that provide clear directional guidance for navigation planning, general description tasks lack explicit objectives to guide the agent’s exploration strategy. Consequently, when directly applying the agent approach to Gemini-2.5-Pro, the model struggles with effective navigation path planning within large regions, as it lacks clear guidance on which critical areas to examine. Since subsequent reasoning heavily depends on the planned navigation path, performance suffers significantly, particularly on datasets like TCGA-BRCA and TCGA-NSCLC.
For example, we observe that on TCGA-NSCLC (Non-Small Cell Lung Cancer) tasks, where all WSIs belong to Non-Small Cell Lung Cancer and the task is to classify subtypes into lung adenocarcinoma and lung squamous cell carcinoma, after applying the agent-based approach, Gemini-2.5-Pro sometimes even misclassifies samples as Small Cell Lung Cancer, demonstrating fundamental diagnostic errors or hallucinations.
This finding validates our progressive training data generation framework, which leverages comprehensive WSI reports to guide region-specific descriptions. These targeted descriptions enable us to prompt Gemini-2.5-Pro to capture multi-scale morphological features across different viewing perspectives. By understanding which anatomical locations contain specific pathological features, we can further prompt Gemini-2.5-Pro to generate higher-quality navigation paths with enhanced spatial awareness, which in turn enables more precise and accurate multi-scale, multi-view reasoning generation. This creates a hierarchical sequence where each stage builds upon the quality of previous stages, resulting in cascading improvements throughout the entire data generation pipeline.
Appendix C Experimental Details
C.1 Details for MIL-based WSI classification
C.1.1 Details for Training and Testing Dataset
To ensure a fair comparison, MIL-based methods are evaluated using identical training and testing datasets such as CPathAgent and other comparative LMMs. All datasets were derived from WSI samples corresponding to the HistGen [57] WSI reports. We select representative subsets suitable for whole slide image classification tasks, comprising the following datasets:
TCGA-RCC (Renal Cell Carcinoma, train: 422, test: 238): A three-class classification task including kidney chromophobe, kidney renal clear cell carcinoma, and kidney renal papillary cell carcinoma.
TCGA-NSCLC (Non-Small Cell Lung Cancer) (train: 642, test: 195): A binary classification task distinguishing between lung adenocarcinoma and lung squamous cell carcinoma.
TCGA-ESCA (Esophageal Carcinoma, train: 83, test: 31): A binary classification task differentiating adenomas and adenocarcinomas from squamous cell neoplasms.
TCGA-BRCA (Breast Invasive Carcinoma, train: 637, test: 226): A binary classification task between invasive ductal carcinoma and invasive lobular carcinoma.
TCGA-THCA (Thyroid Carcinoma, train: 323, test: 119): A three-class classification task encompassing papillary adenocarcinoma, papillary carcinoma columnar cell, and papillary carcinoma follicular variant.
TCGA-BLCA (Bladder Urothelial Carcinoma, train: 227, test: 79): A binary classification task distinguishing papillary transitional cell carcinoma from transitional cell carcinoma.
Additionally, we extract 10% of WSIs from the complete TCGA dataset that are not included in the HistGen dataset, as training the MIL approach requires a separate validation set. Models with the best performance on the validation set are selected for evaluation on the test set.
C.1.2 Details for WSI preprocessing
Our WSI preprocessing pipeline follows the approach established by CPath-omni [10] to implement a multi-scale MIL-based methodology. The preprocessing workflow consists of several key steps: First, we employ CLAM [60] to automatically identify and segment tissue regions by applying appropriate segmentation threshold. From these identified regions, we extract non-overlapping patches of 2048 × 2048 pixels at 40× magnification. To ensure tissue quality, patches are only retained if they contain more than 10% valid tissue area. To capture multi-scale information, each 2048 × 2048 patch undergoes hierarchical subdivision: we maintain the original 2048 × 2048 patch while simultaneously extracting four 1024 × 1024 patches and sixteen 512 × 512 patches from the same region. Features are extracted from all these multi-scale patches using CPath-CLIP [10] and subsequently averaged to generate a unified representation for each 2048-resolution region.
C.1.3 Details for WSI Task-Specific Fine-Tuning
All models are trained for 20 epochs using a fixed learning rate of $1\times 10^{-5}$ without any learning rate scheduling. We employ the Adam optimizer without weight decay, maintaining a batch size of 1 throughout training. To ensure reliable evaluation, we conduct experiments across 5 different random seeds and report the averaged results.
Model Architecture. The MIL framework commonly used for WSI classification includes three learnable components: (1) a fully-connected layer to reduce the dimensionality of features to 256, (2) an attention network to aggregate and transform the instance features, and (3) a final fully-connected layer for making predictions. We experiment with ABMIL [61] and DSMIL [5], both of which share the same fully-connected layers for reducing feature dimensionality and making predictions. For the attention network, ABMIL uses a gated attention mechanism, while DSMIL introduces a dual-stream architecture.
C.2 Hardware and Training Cost of CPathAgent
We employ 8 NVIDIA H800-80G GPUs for CPathAgent training, with computational requirements of 9 hours for Stage 1 (multimodal alignment), 25 hours for Stage 2 (pathology task training), and 39 hours for Stage 3 (agent capability development).
C.3 Training Hyperparameters
For Stages 1 and 2 of CPathAgent training, we follow the hyperparameter configuration of CPath-Omni. For Stage 3 training, we employ 8 NVIDIA H800-80G GPUs and train the model for a single epoch using a per-device batch size of 1 with gradient accumulation of 8 steps. We implement a cosine learning rate scheduler with a base learning rate of 1e-5 and a warmup ratio of 0.03, while fine-tuning the vision tower with a reduced learning rate of 2e-6.
Appendix D Examples and Prompts
D.1 Examples of Region Descriptions Extracted from WSI Report Pairs
As shown in Figure 14, we extract relevant region descriptions from WSI reports by identifying specific diagnostic findings in targeted areas. This process systematically parses pathological reports to highlight key diagnostic information related to the given region, addressing the limitation in TCGA datasets where WSI reports exist but region-specific descriptions are missing.
D.2 Examples of PathMMU-HR² Samples
Figure 15 shows the examples of PathMMU-HR² dataset. These VQA samples demonstrate exceptional quality through the requirement for multi-scale, multi-view reasoning to arrive at accurate diagnoses. As illustrated in the examples, the correct identification of conditions such as fibromatosis (desmoid tumor) and well-demarcated clear cell tumors necessitates careful examination of architectural patterns and cellular features across different magnification levels. The diagnostic process requires the model to analyze tissue structure at lower magnifications while progressively zooming in to evaluate the cellular morphology and stromal characteristics at higher magnifications. This multi-scale analytical approach reflects real-world pathological practice, where definitive diagnoses depend on the integration of macro-architectural features with microscopic cellular details.
D.3 Examples of Region Selection Process Based on WSI Overview
Figure 16 and Figure17 demonstrate the generated region selection results. As illustrated, the model segments the WSI overview into distinct regions and identifies which areas require high-magnification examination. For example, in Figure 16, the model correctly prioritizes core tumor areas for detailed analysis while also identifying critical tumor interface/periphery regions and margin assessment areas to evaluate potential residual tumor presence. In addition, the model appropriately excludes the benign/background breast tissue areas from high-magnification requirements, as these regions primarily consist of adipose tissue and fibrous stroma that appear benign and uninvolved at low magnification, thus optimizing the diagnostic workflow by focusing computational resources on clinically significant regions.
D.4 Examples of Navigation Paths and Multi-Scale Reasoning Generated by CPathAgent
D.4.1 General Navigation Path Planning and Reasoning
Figures 18 and 19 show how CPathAgent navigates and analyzes pathology images like an expert pathologist. For example, as illustrated in Figure 18, CPathAgent begins with a comprehensive overview at low magnification (1.0x), scanning the entire tissue to identify distinct regions and architectural patterns. In this case, it recognizes adipose tissue on the left and dense cellular areas on the right, establishing the overall tissue landscape.
Next, CPathAgent focuses on the main area of concern, examining the primary tumor mass at moderate magnification (2.5x). Here it identifies irregular cell clusters and abnormal tissue architecture that warrant closer investigation. The agent then increases magnification (4.0x) to evaluate cellular details crucial for grading, such as nuclear pleomorphism, mitotic activity, and tubule formation patterns. CPathAgent systematically examines the tumor’s relationship with surrounding tissues, checking the tumor-stroma interface (3.0x) for invasive patterns and examining infiltration into adjacent adipose tissue. It also assesses the immune response by evaluating lymphocytic infiltrates near tumor edges (3.5x) and investigates suspicious ductal structures in other regions (3.0x).
To ensure diagnostic accuracy, CPathAgent performs validation checks by examining additional areas at various magnifications (2.5x, 4.0x), confirming that findings are consistent across different regions of the tissue. This systematic, multi-scale approach allows CPathAgent to integrate architectural observations from low magnification with detailed cellular features from high magnification, ultimately reaching a comprehensive diagnosis of invasive ductal carcinoma with appropriate histological grading—mirroring the thorough diagnostic process used by expert pathologists in clinical practice.
D.4.2 VQA-oriented Navigation Path Planning and Reasoning
Figure 20 demonstrates CPathAgent’s question-guided reasoning approach for VQA task. When presented with a specific question about "morphological patterns describing the relationship between neoplastic cell populations and stromal elements at their interface," CPathAgent tailors its entire analysis strategy to address this inquiry.
CPathAgent begins with a systematic overview at low magnification (1.0x), scanning the tissue to locate regions relevant to the question. It identifies striking heterogeneity across the tissue and recognizes that the lower right quadrant contains the key neoplastic-stromal interface mentioned in the question. With the target area identified, CPathAgent zooms in progressively (2.5x, 4.0x) to characterize the neoplastic component, confirming the presence of nested clear cells with high-grade nuclei. At higher magnification (5.0x), it performs detailed cytological assessment, noting the striking nuclear features that support high-grade classification. CPathAgent then specifically examines the critical interface zone at multiple magnifications (3.5x, 4.0x), systematically documenting the relationship between neoplastic cells and surrounding stromal elements. It identifies key morphological features including the transitional pattern between clear cell neoplasm and eosinophilic tissue, along with dense lymphocytic infiltrate at the junction. Finally, CPathAgent moves to eosinophilic areas (3.0x) to contrast these regions with the neoplastic component, ensuring a comprehensive understanding of the interface characteristics.
This question-driven navigation allows CPathAgent to methodically evaluate multiple-choice options, ultimately selecting the answer that best describes the observed "nested clear cells focally transitioning to solid eosinophilic cells with high-grade nuclei and dense lymphocytes." This demonstrates how CPathAgent’s reasoning process is specifically optimized to answer targeted pathological questions through systematic, multi-scale analysis.
Figure 14: Examples of region description extraction from WSI reports.
Figure 15: Examples of VQAs from the PathMMU-HR² dataset.
Figure 16: An example of region selection based on a WSI overview.
Figure 17: An example of region selection based on a WSI overview.
Figure 18: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure 19: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure 20: An example of CPathAgent’s VQA-oriented navigation path planning and multi-scale, multi-view reasoning process.
D.5 Prompts
This section presents the prompts used in our dataset construction and experiments.
Gemini-2.5-Pro Prompts for Dataset Construction: We design a series of prompts to guide Gemini-2.5-Pro in constructing our pathologist-like planning and reasoning dataset generation pipeline. Figure 21 presents the prompt used to extract region-relevant descriptions from WSI reports, while Figure 22 shows the prompt for region selection using WSI overview images and paired reports. For multi-scale image description generation, Figure 23 displays the prompt that generates detailed descriptions for image patches at different magnifications. Figure 24 contains the prompt for creating navigation plans using coordinate-annotated region overviews and multi-scale descriptions. Finally, Figure 25 shows the prompt that generates pathologist-like diagnostic reasoning based on extracted patches from navigation paths and paired region descriptions.
For VQA dataset construction, we use several specialized prompts: Figure 26 shows the prompt that instructs the model to generate VQA samples, Figure 27 presents the prompt for predicting question answers from text-only inputs (applied to both Gemini-2.5-Pro and DeepSeek-V3), Figure 28 displays the prompt that guides the model to create question-guided navigation plans, and Figure 29 contains the prompt for generating reasoning-based answers.
CPathAgent Training and Inference Prompts: The prompts for CPathAgent follow a similar design, while removing the additional information utilized by Gemini-2.5-Pro (such as pathology reports), focusing on vision-only pathological analysis that replicates clinical diagnostic processes. Figures 30, 31, and 32 present the core prompts used for CPathAgent’s training and inference. These prompts guide CPathAgent to perform navigation planning and reasoning using only visual input. The VQA-specific prompts shown in Figures 33 and 34 display the prompts that used for CPathAgent’s question-guided navigation and reasoning.
Figure 21: Prompt for Gemini-2.5-Pro to extract description related to a given region from the corresponding WSI paired pathology report.
Figure 22: Prompt for Gemini-2.5-Pro to generate region selection results given a WSI overview and its paired pathology report.
Figure 23: Prompt for Gemini-2.5-Pro to generate descriptions for 1 image at 1X scale, 4 images at 2X scale, and 16 images at 4X scale, given these images and their paired region descriptions.
Figure 24: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates and 21 multi-scale image descriptions generated in the previous step.
Figure 25: Prompt for Gemini-2.5-Pro to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path and the overall region description extracted from the previous step.
Figure 26: Prompt for Gemini-2.5-Pro to generate VQA samples given the region overview and the 21 multi-scale image descriptions generated from the previous step.
Figure 27: Prompt for Gemini-2.5-Pro to conduct educated guesses given only the text portion of the generated VQA samples.
Figure 28: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates, 21 multi-scale image descriptions and the question generated in the previous step.
Figure 29: Prompt for Gemini-2.5-Pro to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path, the overall region description extracted from the previous step, and the target question.
Figure 30: Prompt for CPathAgent to generate region selection results given a WSI overview.
Figure 31: Prompt for CPathAgent to generate navigation path planning given the region overview with coordinates.
Figure 32: Prompt for CPathAgent to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path.
Figure 33: Prompt for CPathAgent to generate VQA-oriented navigation path planning given the region overview with coordinates and the question.
Figure 34: Prompt for CPathAgent to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path and the target question.

### v2

@@
-CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic
+envname-P envname#1
 Abstract
 Recent advances in computational pathology have led to the emergence of numerous foundation models.
-However, these approaches fail to replicate the diagnostic process of pathologists, as they either simply rely on general-purpose encoders with multi-instance learning for classification or directly apply multimodal models to generate reports from images.
-A significant limitation is their inability to emulate the diagnostic logic employed by pathologists, who systematically examine slides at low magnification for overview before progressively zooming in on suspicious regions to formulate comprehensive diagnoses.
-To address this gap, we introduce CPathAgent, an innovative agent-based model that mimics pathologists’ reasoning processes by autonomously executing zoom-in/out and navigation operations across pathology images based on observed visual features.
-To achieve this, we develop a multi-stage training strategy unifying patch-level, region-level, and whole-slide capabilities within a single model, which is essential for mimicking pathologists, who require understanding and reasoning capabilities across all three scales.
-This approach generates substantially more detailed and interpretable diagnostic reports compared to existing methods, particularly for huge region understanding.
-Additionally, we construct an expert-validated PathMMU-HR², the first benchmark for huge region analysis, a critical intermediate scale between patches and whole slides, as diagnosticians typically examine several key regions rather than entire slides at once.
-Extensive experiments demonstrate that CPathAgent consistently outperforms existing approaches across three scales of benchmarks, validating the effectiveness of our agent-based diagnostic approach and highlighting a promising direction for the future development of computational pathology.
+These models typically rely on general-purpose encoders with multi-instance learning for whole slide image (WSI) classification or apply multimodal approaches to generate reports directly from images.
+However, these models cannot emulate the diagnostic approach of pathologists, who systematically examine slides at low magnification to obtain an overview before progressively zooming in on suspicious regions to formulate comprehensive diagnoses.
+Instead, existing models directly output final diagnoses without revealing the underlying reasoning process.
+To address this gap, we introduce CPathAgent, an innovative agent-based approach that mimics pathologists’ diagnostic workflow by autonomously navigating across WSI through zoom-in/out and move operations based on observed visual features, thereby generating substantially more transparent and interpretable diagnostic summaries.
+To achieve this, we develop a multi-stage training strategy that unifies patch-level, region-level, and WSI-level capabilities within a single model, which is essential for replicating how pathologists understand and reason across diverse image scales.
+Additionally, we construct PathMMU-HR², the first expert-validated benchmark for large region analysis.
+This represents a critical intermediate scale between patches and whole slides, reflecting a key clinical reality where pathologists typically examine several key large regions rather than entire slides at once.
+Extensive experiments demonstrate that CPathAgent consistently outperforms existing approaches across benchmarks at three different image scales, validating the effectiveness of our agent-based diagnostic approach and highlighting a promising direction for computational pathology.
 1 Introduction
@@
 Pathology serves as the gold standard for diagnosing numerous diseases, particularly cancer.
-Advances in digital pathology have enabled the digitization and computational analysis of histopathological slides, creating promising opportunities for AI-assisted diagnostics.
-Recently, various pathology foundation models have emerged to meet the growing demand for automated analysis in this field.
-Current foundation models in computational pathology primarily fall into two categories (as shown in the left part of 1): encoder-based models like DINO[1; 2] or CLIP [3] that extract patch-level representations from Whole Slide Images (WSIs) followed by Multi-Instance Learning (MIL) [4; 5; 6] aggregation to perform diagnosis, and large multimodal models (LMMs) that process both visual data and text instructions to generate pathology reports or other outputs [7; 8; 9; 10].
-Despite their impressive benchmark performance, these approaches share a critical limitation: they fail to incorporate the intermediate reasoning processes that expert pathologists employ, limiting their interpretability and clinical applicability.
-In clinical practice, pathologists employ a methodical approach to slide examination, starting with low-magnification scanning (typically 2-4×) to assess tissue architecture and overall morphology.
-They then strategically navigate to regions of interest showing atypical features or diagnostic importance, progressively increasing magnification (10×, 20×, and 40×) to evaluate fine-grained histopathological features.
-Throughout this process, pathologists continuously shift between different magnifications and fields of view, integrating spatial and morphological information across multiple scales to formulate a comprehensive diagnosis.
-Current computational models bypass this sophisticated visual reasoning process, resulting in: (1) insufficient interpretability, as they cannot demonstrate visual navigation and multi-scale reasoning; (2) vulnerability to hallucinations and shortcut learning, potentially fixating on spurious correlations rather than relevant histological patterns; and (3) poor alignment with clinical workflows, undermining their credibility in diagnostic settings.
-Drawing inspiration from recent breakthroughs in agent-based systems [11; 12; 13], which have demonstrated remarkable capabilities across diverse domains including strategic gameplay [14], web navigation [15], complex task automation [16], and multimodal content generation [17], we introduce CPathAgent, a novel framework that emulates the adaptive visual reasoning of expert pathologists.
-Unlike existing computational pathology models that operate as black-box classifiers or descriptors, CPathAgent implements a dynamic, perception-driven approach that mirrors pathologists’ hierarchical examination strategy in clinical practice.
-CPathAgent performs strategic operations: zooming in to examine cellular details, zooming out to maintain contextual awareness, and navigating across regions of interest, mimicking how pathologists scan slides at low magnification to identify suspicious areas before zooming in to examine fine-grained histological patterns before making diagnosis.
-This represents a fundamental paradigm shift in computational pathology from static, outcome-focused models to interactive, reasoning-based systems with greater clinical utility.
-Specifically, CPathAgent mimics pathologists’ diagnostic workflows in a multi-stage process, as shown in the right part of Figure 1: (1) Global screening and region selection: CPathAgent initially analyzes a low-resolution WSI view to identify regions of interest (16000×16000 region at 40×) exhibiting suspicious histological features for detailed examination.
-(2) Navigation planning: For each prioritized region, CPathAgent plans a navigation path defining a sequence of view coordinates at varying magnifications, determining when to pan, zoom in, or zoom out based on the observed features within the huge region.
-(3) Multi-scale multi-view sequence reasoning: CPathAgent executes the planned navigation path by analyzing specific fields of view at different magnifications, then synthesizes observations across these views to perform step-by-step diagnostic reasoning.
-Our work makes four significant contributions to advance agent-based computational pathology:
-1) We introduce CPathAgent, which is a agent-based framework that emulates a pathologist’s diagnostic workflow through dynamic navigation planing and multi-magnification multi-view analysis, enabling interpretable, reasoning-based assessments.
-2) We develop PathMMU-HR², the first visual question answering benchmark specifically designed for high-resolution huge pathology image regions (16000×16000 pixels), rigorously validated by three professional pathologists to ensure clinical relevance and diagnostic accuracy.
-3) We introduce a multi-stage training strategy: An approach that equips CPathAgent with cognitive and reasoning capabilities across different scales, from patch-level details to whole-slide contexts, within a unified diagnostic framework.
-4) We conduct extensive evaluations across benchmarks at multiple scales, ranging from patches to large regions to WSIs, highlighting CPathAgent’s strong performance and potential clinical utility.
+The advent of digital pathology has enabled the digitization and computational analysis of histopathological slides, opening transformative opportunities for AI-assisted diagnostics.
+Recently, numerous pathology foundation models have emerged to automate and enhance diagnostic workflows.
+These foundation models (Figure 1, left) adopt two primary architectural paradigms.
+The first employs encoder-based models such as DINO[1; 2] or CLIP [3] to extract patch-level representations from WSIs, which are then aggregated via Multi-Instance Learning (MIL) [4; 5; 6] for diagnostic predictions.
+The second utilizes large multimodal models (LMMs) that jointly process visual data and text instructions to generate pathology reports or structured outputs [7; 8; 9; 10].
+While these approaches achieve impressive benchmark performance, they share a fundamental mismatch with clinical practice.
+In real-world diagnostics, pathologists employ a systematic multi-scale examination strategy.
+They begin by scanning slides at low magnification (2-4×) to assess overall tissue architecture and identify regions of interest, then strategically navigating to atypical regions while progressively increasing magnification (10×, 20×, 40×) to examine fine-grained details.
+Throughout this process, pathologists continuously integrate observations across scales while conducting step-by-step diagnostic reasoning.
+However, current models bypass this systematic sequential process by directly aggregating information across the WSI or processing visual inputs in a single forward pass, leading to two critical limitations.
+First, weak supervision: relying solely on slide-level labels without explicit guidance on diagnostically important regions or when to examine specific magnifications makes them vulnerable to shortcut learning [6], exploiting spurious correlations (staining artifacts, background patterns) rather than genuine pathological features.
+Second, lack of interpretability and verifiability: producing only final predictions without providing natural language descriptions of the intermediate reasoning prevents pathologists from understanding what the AI actually discovered and validating whether its findings are based on clinically relevant features.
+To address these limitations, we introduce CPathAgent, a novel agent-based framework inspired by recent advances in AI agent systems that have achieved remarkable success from strategic gameplay [11] and web navigation [12] to task automation [13] and multimodal content generation [14; 15; 16; 17].
+The key to their success lies in decomposing complex tasks into sequential reasoning steps with dynamic decision-making.
+Following this paradigm, CPathAgent bridges the gap between clinical practice and AI-assisted diagnosis by rethinking pathology diagnosis as adaptive visual reasoning.
+Unlike existing black-box models that treat diagnosis as single-shot classification, CPathAgent emulates pathologists’ examination strategy through WSI navigation with dynamic magnification adjustment and multi-view synthesis, conducting explicit step-by-step reasoning through natural language to enable interpretable and verifiable diagnostic assessments.
+Specifically, CPathAgent performs a three-stage diagnostic workflow (Figure 1).
+In the global screening stage, CPathAgent analyzes a low-resolution WSI overview to identify suspicious regions that warrant detailed examination.
+In the navigation planning stage, CPathAgent plans an exploration strategy for each selected region, determining a sequence of view coordinates and magnification levels to investigate, mimicking how pathologists mentally map their examination before diving into details.
+Finally, in the multi-scale reasoning stage, the planned navigation path is executed through systematic analysis of fields of view at different magnifications (e.g., zooming in for cellular details, moving laterally across areas, zooming out for context), and synthesizes observations via step-by-step diagnostic reasoning to reach a conclusion.
+Overall, this work makes the following contributions:
+1) CPathAgent framework: We propose an agent-based diagnostic system that performs diagnosis through strategic WSI navigation and multi-scale visual reasoning, mirroring pathologists’ systematic examination process while providing interpretable step-by-step rationales.
+2) PathMMU-HR² benchmark: We develop the first visual question answering benchmark specifically designed for high-resolution pathology image regions (16000×16000 pixels), rigorously validated by three professional pathologists to ensure clinical relevance and diagnostic accuracy.
+3) Multi-stage training strategy: We design a progressive training approach that equips CPathAgent with perceptual and reasoning capabilities across different image scales within a unified framework.
+4) Comprehensive evaluation: We conduct experiments at multiple scales, from patches and large regions to WSIs, highlighting CPathAgent’s strong performance and practical clinical utility.
 2 Related Work
-2.1 Pathology Language-Vision Models
-Recent years have witnessed substantial progress in developing foundation models specifically tailored for computational pathology.
-These models broadly fall into two categories: encoder-based models and large multimodal models (LMMs).
-Encoder-based models primarily aim to extract patch-level representations from whole-slide images (WSIs).
-Early efforts in this area relied on ImageNet-pretrained convolutional neural networks (CNNs) [18; 19] as feature extractors, combined with multi-instance learning (MIL) techniques [4; 5; 6] to aggregate these features for slide-level prediction.
+2.1 Pathology Foundation Models
+Recent years have seen significant progress in developing foundation models for computational pathology, broadly categorized into encoder-based models and large multimodal models (LMMs).
+Encoder-based models focus on extracting rich patch-level representations from WSIs.
+Early efforts in this area relied on ImageNet-pretrained CNNs [18; 19] as feature extractors, combined with multi-instance learning (MIL) [4; 5; 6] to aggregate these features for slide-level prediction.
 More recently, the field has shifted toward pathology-specific foundation models that adopt self-supervised learning paradigms.
 Architectures such as DINO [1; 2] and CLIP [3], when trained on large-scale pathology datasets, have demonstrated superior feature extraction capabilities [20; 8; 21; 22; 23].
-These improved representations, integrated via MIL or incorporated into pretrained WSI transformers, have significantly advanced the state of the art in WSI-level classification.
-Large multimodal models (LMMs) for pathology have emerged in parallel, enabling the processing of pathology images alongside text-based instructions to generate textual outputs.
+These improved representations, integrated via MIL or incorporated into pretrained WSI transformers [20; 9], have significantly advanced the WSI-level classification performance.
+LMMs for pathology have emerged in parallel, enabling joint processing of pathology images and text instructions to generate textual diagnostic outputs.
 These models integrate visual and linguistic information through various designs.
-Patch-level MLLMs [7; 24; 8; 25; 26] adapt general vision-language frameworks like LLaVA [27] to interpret pathology image patches, while WSI-level models [9; 28; 29] aim to process entire whole-slide images for generating reports or answering diagnostic questions.
+Patch-level LMMs [7; 24; 8; 25; 26] adapt general vision-language frameworks like LLaVA [27] for patch-level pathology image interpretation, while WSI-level models [9; 28; 29] aim to process entire WSIs for generating reports or answering diagnostic questions.
 Recent unified frameworks [10] attempt to handle both patches and WSIs within a single model architecture.
-Despite these advances, current models typically process pathology inputs in a static and fixed manner, which contrasts with the dynamic, interactive diagnostic workflow of expert pathologists, who routinely navigate across regions of interest and examine tissues at multiple magnification levels, integrating contextual information over time to formulate a diagnosis.
+Despite these advances, current models typically process pathology inputs in a static manner, contrasting with expert pathologists who dynamically navigate regions of interest, examine tissues at multiple magnifications, and integrate contextual information to formulate diagnoses.
 2.2 Agent-based Models
 Agent-based models represent an emerging paradigm in AI research, characterized by systems that can perceive their environment, make decisions, and take actions to achieve specific goals [30].
-Recent advances in large language models (LLMs) have catalyzed significant progress in this domain, enabling more sophisticated reasoning and planning capabilities.Notable examples include WebGPT [15] and WebShop [31], which navigate web environments to complete information-seeking and shopping tasks; ReAct [32], which interleaves reasoning and action steps for improved task performance; AutoGPT [33] and AutoGen [34], which decompose complex tasks into manageable subtasks with minimal human supervision; and task-specific agents for code development [17], creative content generation [35], scientific research [36], and clinical decision-making and discovery that simulate physician workflows in hospital environments [37; 38; 39; 40].
-These agent models typically incorporate several key components [11]: perception modules to process environmental inputs, reasoning modules to interpret observations and plan actions, and execution modules to implement decided actions.
+Recent advances in large language models (LLMs) have catalyzed significant progress in this domain, enabling more sophisticated reasoning and planning capabilities.
+Notable examples include WebGPT [12] and WebShop [31], which navigate web environments to complete information-seeking and shopping tasks; ReAct [32], which interleaves reasoning and action steps for improved task performance; AutoGPT [33] and AutoGen [34], which decompose complex tasks into manageable subtasks with minimal human supervision; and task-specific agents for code development [14], creative content generation [35], scientific research [36], and clinical decision-making and discovery that simulate physician workflows in hospital environments [37; 38; 39; 40].
+These agent models typically incorporate several key components [15]: perception modules to process environmental inputs, reasoning modules to interpret observations and plan actions, and execution modules to implement planned actions.
 This architecture enables agents to adapt to dynamic environments and accomplish complex, multi-step tasks through iterative observation-reasoning-action cycles.
-Although agent-based approaches have demonstrated success across various domains, their application to computational pathology remains largely unexplored.
-PathFinder [41] represents a notable early attempt at introducing agent-based systems for WSI analysis.
-The system segments WSIs into fixed 512×512 patches and employs a U-Net-based architecture to generate importance maps.
-To be specific, PathFinder feeds WSI thumbnails to generate initial importance maps, selects high-probability regions for description generation, and regenerate subsequent importance maps based on previous descriptions and thumbnails.
-While PathFinder introduces agent-based concepts to WSI analysis, it primarily operates through region selection at a fixed magnification level within predefined 512×512 patches.
-Pathologists, however, employ a more flexible diagnostic workflow with dynamic multi-scale navigation—beginning with low-magnification overviews and continuously adjusting zoom levels while moving across tissue regions based on morphological observations to integrate information across different scales.
-To address this gap, our CPathAgent takes a step further by incorporating multi-scale navigation capabilities that mimic pathologist behavior.
+Despite success in various domains, agent-based approaches remain largely unexplored in computational pathology.
+PathFinder [41] represents a successful early attempt at introducing agent-based systems for WSI analysis.
+The system segments WSIs into fixed 512×512 patches and employs a multi-agent framework that iteratively generates importance maps to select diagnostically relevant regions, produces natural language descriptions of these patches, and then regenerates subsequent importance maps conditioned on previous descriptions.
+However, PathFinder primarily operates through region selection at a fixed magnification level within predefined 512×512 patches, lacking the dynamic multi-scale navigation capability that characterizes pathologist workflows.
+To address this gap, CPathAgent takes a step further by incorporating multi-scale navigation capabilities that mimic pathologist behavior.
 It autonomously decides where to examine, when to zoom in or out, and how to navigate across WSIs, closely mimicking the iterative diagnostic workflow of expert pathologists.
-Notably, CPathAgent provides real-time verbalization of its diagnostic reasoning, effectively simulating the internal thought processes that pathologists experience during slide examination (specific examples can be found in Appendix Figure 16 to 20).
+Notably, CPathAgent provides verbalization of its diagnostic reasoning, effectively simulating the internal thought processes that pathologists experience during slide examination (specific examples can be found in Appendix Figures A12 to A16).
 This approach enables enhanced interpretability through traceable reasoning paths and improved diagnostic performance.
 2.3 Pathology Benchmarks and Datasets
-Benchmark datasets play a crucial role in advancing computational pathology by providing standardized evaluation frameworks.
-Traditional whole-slide image (WSI) benchmarks like Camelyon16 [42] and TCGA [43] focused primarily on slide-level classification tasks.
-More recently, datasets like WSICaption[44] and WSI-VQA[45] have emerged, targeting whole-slide image description generation and question-answering tasks across entire slide images, respectively.
+Benchmark datasets provide standardized evaluation frameworks crucial for advancing computational pathology.
+Traditional WSI benchmarks like Camelyon16 [42] and TCGA [43] focused primarily on slide-level classification tasks.
+More recently, datasets like WSICaption[44] and WSI-VQA[45] have expanded the scope to more complex tasks, including WSI description generation and VQA.
 In parallel, several patch-level benchmarks have emerged for evaluating multimodal models in pathology.
-PathVQA [46] introduced the first visual question answering (VQA) tasks for pathology images, drawing on textbook image -caption pairs.
-QuiltVQA expanded this approach by extracting image-caption pairs from pathology lectures on YouTube, resulting in higher-quality QA pairs due to the educational nature of these materials.
-PathMMU [47] marks a more recent advance, offering a comprehensive benchmark across diverse pathology sources.
-Its key strength is the expert validation process, resulting in a professionally curated dataset that better reflects clinical practice.
-Despite recent progress, a critical gap remains in existing benchmarks: they focus primarily on either small, localized patch-level tasks or broad whole-slide analysis, overlooking an essential intermediate scale, the huge region, which plays a central role in real-world diagnostic reasoning.
-In routine practice, pathologists typically begin by scanning the whole slide images to identify suspicious areas, and then concentrate their diagnostic attention on large, yet manageable regions containing the relevant pathology.
+PathVQA [46] introduced the first VQA tasks for pathology images, drawing on textbook image-caption pairs.
+QuiltVQA [47] expanded this approach by extracting image-caption pairs from pathology lectures on YouTube, resulting in higher-quality VQA pairs due to the educational nature of these materials.
+PathMMU [48] marks a more recent advance, offering a comprehensive benchmark across diverse pathology sources validated by expert pathologists to ensure clinical relevance.
+Despite this progress, existing benchmarks predominantly focus on either localized patch-level tasks or broad WSI analysis, overlooking a critical intermediate scale: the huge region.
+This scale is particularly important because it mirrors real-world diagnostic workflows.
+In routine practice, pathologists typically begin by scanning the WSIs to identify suspicious areas, then focus their diagnostic attention on these large, yet manageable regions.
 It is often within these huge regions that key diagnostic decisions are made, as they provide sufficient context for interpretation.
 To address this limitation, we introduce PathMMU-HR², a benchmark specifically designed for huge region analysis in pathology.
-By targeting this intermediate scale—situated between isolated patches and WSIs, PathMMU-HR² better aligns with the way pathologists actually examine and interpret tissue.
+By targeting this intermediate scale between isolated patches and WSIs, PathMMU-HR² better aligns with how pathologists actually examine and interpret tissue.
 It provides a more controlled, context-rich, and diagnostically meaningful setting for evaluating multimodal models, especially those incorporating agent-based reasoning across different scales.
@@
 Figure 2: Illustration of CPathAgent framework, which mimics the diagnostic workflow of pathologists via global screening, navigation planning, and multi-scale reasoning across sequential views.
-In this section, we provide a comprehensive description of our proposed approach.
-We detail the construction of CPathAgent-Instruct, our instruction-tuning dataset used to train CPathAgent, the development of the PathMMU-HR2 benchmark dataset, and the specific training methodology implemented for the CPathAgent model.
-3.1 Overview of CPathAgent Reasoning Workflow
+In this section, we detail the construction of CPathAgent-Instruct training data, the PathMMU-HR² benchmark, and the CPathAgent model architecture with its multi-stage training process.
+3.1 Reasoning Workflow of CPathAgent
 To simulate expert-level diagnostic reasoning, we design CPathAgent as an agent-based system capable of dynamic region selection, strategic navigation planning, and multi-scale reasoning.
 As illustrated in Figure 2, the entire diagnostic process is decomposed into three key stages:
-1) Global Screening: CPathAgent first scans the low-resolution thumbnail of a whole-slide image to identify and cluster suspicious regions.
-Uninformative areas are filtered out, while large diagnostically relevant regions (16000$\times$16000 pixels) are preserved for further analysis.
-2) Navigation Planning: For each preserved huge region, the model plans a navigation path across spatial coordinates and magnification levels (e.g., 1$\times$, 3$\times$, 4.5$\times$).
-The goal is to plan sequentially examine areas that are most likely to yield diagnostic insights.
-3) Multi-scale Multi-view Sequence Reasoning: Along the planned navigation path, the model observes each field of view, integrates visual features across scales and generates step-by-step reasoning descriptions similar to a pathologist’s thought process explaining the purpose of actions like zooming in and moving and finally synthesizes findings into a coherent diagnostic conclusion.
+1) Global Screening: To address the computational challenge of gigapixel WSIs, CPathAgent employs a coarse-to-fine region selection that mirrors pathologists’ systematic screening protocol.
+[…]
+2) Navigation Planning: For each preserved huge region $r_{i}$, the model generates a dynamic navigation plan as a sequence of viewing steps.
+Each step is specified by a tuple $(x,y,m,o)$, where $(x,y)$ denotes spatial coordinates in normalized space $[0,1]$, $m$ represents the magnification level relative to the original huge region (where $1.0\times$ corresponds to the full view), and $o$ describes the diagnostic focus for that position (i.e., what features need to be observed).
+The navigation plan $\mathcal{P}_{i}=\{(x_{1},y_{1},m_{1},o_{1}),\ldots,(x_{T},y_{T},m_{T},o_{T})\}$ is generated autoregressively, where each step is conditioned on the visual content of the region and WSI source information (e.g., lung, colon, etc.).
+3) Multi-scale Multi-view Sequence Reasoning: Following the planned navigation path $\mathcal{P}_{i}$, the model receives the complete sequence of cropped images corresponding to all viewing steps $\{(x_{1},y_{1},m_{1},o_{1}),\ldots,(x_{T},y_{T},m_{T},o_{T})\}$ as input at once.
+Given this multi-view image sequence along with their associated diagnostic focus descriptions and spatial coordinates, the model performs holistic reasoning across the entire viewing trajectory in a first-person perspective similar to a pathologist’s examination process.
+The reasoning process systematically observes features across multiple scales and positions, establishes and refines diagnostic hypotheses by cross-referencing evidence, and maintains logical continuity throughout the navigation sequence.
+Finally, the model synthesizes all observations into a coherent pathological report that summarizes the diagnostic findings.
+Illustrative examples, including global screening results, navigation path planning, and multi-scale reasoning outputs, are shown in Appendix Figure A12 to Figure A15.
+The prompts used for CPathAgent through these stages are provided in Appendix Figure A26 to Figure A28.
 3.2 CPathAgent-Instruct Dataset Construction
-To effectively train our agent-based model to simulate the complex diagnostic reasoning of expert pathologists, we construct the CPathAgent-Instruct dataset, which consists of several subsets targeting different critical reasoning capabilities across multiple scales.
-Unlike approaches that rely solely on prompt engineering during inference, our focus is on generating high-quality instruction data to train open-source models capable of expert-level pathology reasoning.
-The data generation strictly follows the reasoning workflow illustrated in Figure 1 but leverages a additional information to enhance data quality, summarized as follows:
-Region Selection Dataset: We start by downscaling each WSI by a factor of 32 to create navigable thumbnails.
-We define huge regions (16000×16000 pixels at 40× magnification) with 5% overlap between adjacent regions to preserve continuity across region boundaries.
-We prompt Gemini-2.5-Pro with carefully crafted prompts and WSI paired reports (detailed in Appendix D) to generate structured region annotations that include: (1) region groupings with descriptive names, (2) prioritized ID lists representing recommended viewing order, and (3) flags indicating necessity for high-magnification examination.
-This process produces thumbnail-region selection pairs that emulate a pathologist’s systematic WSI screening strategy.
-Viewing Path Planning Dataset: To teach the model strategic examination sequences across multiple magnifications, we implement a three-step reference-guided approach: (1) We extract region-specific content from expert-written WSI reports by prompting Gemini-2.5-Pro with specific region and WSI report (upper left panel of Figure 3); (2) We first crop patches at three magnification levels (1× for comprehensive overview, 2× for four quadrant views, and 4× for sixteen high-resolution details), then send both these multi-scale patches and the previously extracted region-specific captions to Gemini-2.5-Pro, this guidance from expert descriptions enables the model to generate more accurate and diagnostically relevant patch descriptions (upper right panel of Figure 3); Having detailed descriptions at different scales guides the model to more effectively generate navigation planning; (3) We then annotate images with coordinate grids (0.1 relative position intervals) as reference, and use the previously generated patch descriptions to prompt Gemini-2.5-Pro to produce viewing paths.
-We specifically require generation of variable-scale paths that can navigate to any scale and position, rather than fixed-scale routes.
-These steps produce paired huge region-navigation planned path data that forms the foundation of our viewing path dataset.
-Multi-Image Understanding and Reasoning Dataset: We extract image crops at specified magnification levels along the generated viewing paths (lower right panel of Figure 3).
-We pair these multi-scale image sets with WSI report fragments and prompt Gemini-2.5-Pro to generate step-by-step visual reasoning sequences.
-The resulting dataset contains structured reasoning chains that integrate observations across magnification levels, identify specific morphological features, and formulate diagnostic interpretations based on collective evidence, enabling CPathAgent to develop sophisticated multi-scale reasoning capabilities.
+In contrast to methods relying on inference-time prompting with closed-source models (e.g., OpenAI o3, Gemini-2.5-Pro), we focus on curating high-quality data to train open-source models for expert-level agent-based pathology analysis.
+To this end, we construct the CPathAgent-Instruct dataset, comprising corresponding subsets that target the three critical stages detailed in Section 3.1.
+While we use Gemini-2.5-Pro following the approaches in Section 3.1 to synthesize training data, prompting alone proves insufficient for expert-level performance.
+We therefore incorporate reference information such as WSI reports as guidance to ensure data quality.
+The construction process is as follows:
+Source Data: We use WSI reports from HistGen [49] and corresponding WSIs from TCGA, splitting 80% (5,254 WSIs) for training and 20% for testing and PathMMU-HR² generation.
+We apply patient-level splitting to ensure no data leakage across sets.
+Details are provided in Appendix A.
+Global Screening Subset: This dataset follows the same WSI processing as CPathAgent’s global screening stage, with one key difference: we leverage Gemini-2.5-Pro, using WSI overview as primary inputs and corresponding WSI reports as additional guidance, to generate high-quality structured outputs (region groupings, priority scores, and examination flags) as training data.
+[…]
+Multi-scale Multi-view Sequence Reasoning Subset: The synthesis of this subset follows the same process as CPathAgent’s multi-scale reasoning stage (Section 3.1).
+In addition to the cropped image sequences extracted along the planned navigation paths, we provide the region description $D_{\text{region}}$ as reference guidance to Gemini-2.5-Pro for generating high-quality step-by-step reasoning chains.
 Figure 3: Overview of the generation process for CPathAgent’s navigation planning subset and multi-scale multi-view sequence reasoning instruction-tuning data.
-VQA-oriented Dataset: To enable CPathAgent to generate question-oriented dynamic viewing paths, we develop a specialized dataset for huge region VQA.
-First, we generate diverse clinical questions based on previously generated multi-scale patch descriptions.
-Next, we adapt our viewing path planning approach by prompting Gemini-2.5-Pro with these questions alongside region descriptions to generate examination paths optimized for answering each clinical query.
-Finally, we synthesize step-by-step reasoning chains that navigate through relevant visual evidence to reach clinically sound answers.
-Unlike our previous datasets guided by comprehensive region descriptions, this dataset is driven by specific visual questions, training CPathAgent to adapt its examination strategy based on question context—selecting optimal magnification levels and fields of view to efficiently answer diagnostic queries.
-Overall, we generated 278K instruction-tuning data points that collectively capture the dynamic, multi-scale reasoning processes that mimic expert pathologists’ diagnostic workflows.
-The resulting CPathAgent-Instruct dataset equips our model with advanced capabilities essential for computational pathology: strategic region selection, examination planning, cross-magnification feature integration, and targeted diagnostic reasoning and question answering.
-For more detailed dataset statistical information, please kindly refer to Appendix A.
-3.3 PathMMU-HR2 Dataset Construction
-To validate CPathAgent’s capabilities in huge region analysis, we developed the PathMMU-HR2 benchmark dataset.
-We ensured diverse pathological representation by sampling regions across various tissue types (e.g., TCGA-BRCA, TCGA-LUAD, etc.) from The Cancer Genome Atlas (TCGA), and selecting ultra-regions from different histological groups (as detailed in section 3.2) to capture a wide range of pathologically significant structures.
-Using an approach similar to CPathAgent-Instruct development, we generated VQA pairs that specifically require integration of features across multiple magnification levels.
-Three board-certified pathologists rigorously evaluated all generated VQA pairs.
-Their assessment focused on the clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration to answer questions, and alignment with standard pathology practice.
-VQA pairs that failed to meet any of the established quality criteria were eliminated from the dataset.
-The finalized PathMMU-HR2 benchmark contains 1,688 expert-validated VQA instances specifically designed to evaluate multi-scale reasoning capabilities in huge regions.
-Figure 4: Overview of architecture and three-stage training pipeline of CPathAgent.
+VQA-oriented Subset: To enable CPathAgent to handle both general diagnosis and VQA scenarios, this is an additional subset that adapts the navigation planning process to question-driven contexts.
+We first prompt Gemini-2.5-Pro to generate pathology questions from multi-scale patch descriptions $\{d_{1},d_{2},\ldots,d_{21}\}$, then generate question-oriented navigation paths by prompting with both the question and region description $D_{\text{region}}$.
+Finally, we prompt Gemini-2.5-Pro with the cropped image sequences along with the question and region description $D_{\text{region}}$ to produce question-oriented reasoning chains.
+This trains CPathAgent to dynamically adjust its examination strategy based on question-oriented targeted areas and magnifications to efficiently answer diagnostic questions.
+Overall, CPathAgent-Instruct comprises 278K instruction-tuning samples for training CPathAgent’s pathologist-like agent capabilities. Detailed data statistics are provided in Appendix A.
+3.3 PathMMU-HR² Dataset Construction
+To evaluate CPathAgent’s multi-scale reasoning capabilities on huge regions, we construct PathMMU-HR², a Huge Region Huge Resolution benchmark specifically designed to assess the model’s capability for huge region analysis.
+We sample huge regions from the held-out TCGA test set, ensuring broad coverage across tissue types (e.g., TCGA-BRCA, TCGA-LUAD).
+Following the construction pipeline of CPathAgent-Instruct VQA-oriented subset, we generate VQA pairs that require synthesizing observations across different scales, necessitating multi-scale reasoning for accurate diagnosis.
+Three board-certified pathologists independently review and filter generated VQA pairs based on clinical relevance and diagnostic accuracy, necessity of multi-scale integration, and alignment with standard pathology practice.
+The final PathMMU-HR² comprises 1,668 expert-validated VQA samples, providing a robust benchmark for evaluating huge region analysis capabilities.
 3.4 CPathAgent Model Architecture and Training
-CPathAgent employs a multimodal architecture built upon the LLaVA-OneVision [48] framework, with significant adaptations tailored for computational pathology tasks.
-At its core, CPathAgent integrates two powerful components: Qwen3-14B [49], a powerful LLM that provides strong reasoning and generation capabilities, and CPath-CLIP, a powerful pathology-specific CLIP model trained on over 700,000 high-quality pathology image-caption pairs.
-These components are connected through a two-layer MLP projection layer.
-We implement a multi-stage training strategy, as shown in Figure 4 which is a progressive three-stage approach that systematically builds CPathAgent’s capabilities from foundational multimodal alignment to advanced agent-based diagnostic reasoning:
-Stage 1 focuses exclusively on aligning the feature spaces between CPath-CLIP and the Qwen3-14B LLM.
-Only the two-layer MLP projection layer connecting the vision encoder to the language model is trained, while keeping both base models frozen.
-This training utilizes the CPath-PatchCaption dataset [10], enabling effective pre-alignment of visual and language representations without disrupting the pre-trained capabilities of either component.
-Stage 2 builds upon the aligned feature spaces from Stage 1, where we train CPathAgent on comprehensive pathology tasks using the CPath-PathInstruct dataset [10].
-During this stage, all model parameters are unfrozen for fine-tuning, allowing CPathAgent to master diverse patch-level tasks such as visual question answering, image classification, captioning, and diagnostic reasoning.
-This stage establishes the foundation for CPathAgent’s higher-level diagnostic capabilities.
-Stage 3, the most crucial phase, focuses on transforming CPathAgent from a multimodal model into an agent capable of emulating pathologists’ diagnostic workflows.
-We combined our specialized CPathAgent-Instruct dataset with 20% of CPath-Instruct for training.
-This integration equips CPathAgent with agent capabilities that mimic pathologists’ whole-slide image navigation and large region reasoning processes, while preserving the strong patch-level understanding developed in previous stages.
-As in Stage 2, all model parameters remain unfrozen during this stage.
-Our multi-stage training strategy produces a versatile model that excels at both detailed patch-level analysis and strategic large-region and whole-slide image navigation and examination, making CPathAgent a powerful tool that can adapt to various clinical scenarios requiring different levels of pathological analysis.
+CPathAgent builds upon the LLaVA-OneVision [50], integrating Qwen3-14B [51] as the LLM backbone and CPath-CLIP [10] as the vision encoder, connected via a two-layer MLP.
+We adopt a three-stage progressive training strategy that systematically develops CPathAgent’s capabilities, advancing from foundational multimodal understanding to agent-based diagnostic reasoning.
+To be specific, stages 1 and 2 follow the CPath-Omni [10] training protocol: stage 1 aligns vision and language components using CPath-PatchCaption with only the MLP trainable, while stage 2 fine-tunes all parameters on CPath-PathInstruct to develop comprehensive patch-level pathology understanding, including VQA, classification, and image description capabilities.
+Notably, stage 3 is the critical phase that transforms CPathAgent into a pathologist-like agent.
+During this stage, we train on our specialized CPathAgent-Instruct dataset, supplemented with 20% of CPath-Instruct data, with all parameters unfrozen.
+This phase equips CPathAgent with advanced agent capabilities that emulate pathologists’ diagnostic workflows while preserving the strong patch-level analysis skills from earlier stages.
+This multi-stage approach yields a versatile model that seamlessly integrates fine-grained patch-level analysis with pathologist-like agent-based WSI navigation.
 4 Experiments
@@
 4.1 Patch Understanding Evaluation
-Since patch-level understanding serves as the foundation for reasoning over larger regions and for planning across WSIs, we first evaluate CPathAgent’s ability to interpret standard-resolution pathology image patches.
-We conduct this evaluation on PathMMU, currently the largest and most diverse expert-verified pathology dataset.
-We compare CPathAgent with the most advanced general-purpose LMMs, including InstructBLIP-FLAN-T5 XXL[50], LLaVA-1.5-13B[51], LLaVA-OneVision[48] series, Qwen2.5-VL[52] series, GPT-4V[53], GPT-4.1[54] series and Gemini-2.5[55] series, along with domain-specific pathology LMMs such as LLaVA-Med[56], Quilt-LLaVA[8], PathGen-LLaVA[26] and CPath-Omni[10].
-Results: CPathAgent significantly outperforms both general-purpose and pathology-specific LMMs across all subsets.
-As shown in Table 1, it achieves the highest performance across all subsets, with 80.5% on the tiny test set and 78.6% on the full test set.
-This shows a clear margin over the strongest general-purpose model (Gemini-2.5-Pro, 68.7% Tiny / 67.5% All) and surpasses the SOTA domain-specific model (CPath-Omni, 72.4% Tiny / 72.2% All).
-While part of CPathAgent’s patch-level performance can be attributed to the training data inherited from CPath-Omni, CPathAgent benefits further from the third-stage training with multi-scale and reasoning-focused data.
-This helps it develop stronger reasoning ability and builds a solid foundation for simulating expert-like behaviors such as zooming, navigation, and diagnostic thinking across WSIs.
+Since patch-level understanding serves as the foundation for reasoning over larger-region reasoning, we first evaluate CPathAgent’s ability to interpret standard-resolution pathology patches on PathMMU [48], currently the largest and most diverse expert-verified pathology dataset.
+We compare CPathAgent with the most advanced general-purpose LMMs, including InstructBLIP-FLAN-T5-XXL[52], LLaVA-1.5-13B[53], LLaVA-OneVision[50] series, Qwen2.5-VL[54] series, GPT-4V[55], GPT-4.1[56] series and Gemini-2.5[57] series, along with domain-specific pathology LMMs such as LLaVA-Med[58], Quilt-LLaVA[8], PathGen-LLaVA[26] and CPath-Omni[10].
+Results: CPathAgent significantly outperforms both general-purpose and pathology-specific LMMs.
+As shown in Table 1, it achieves the highest performance across all subsets, with 80.5% on the tiny test set and 78.6% on the full test set, surpassing the strongest general-purpose model (Gemini-2.5-Pro, 68.7% Tiny / 67.5% All) and surpasses the SOTA domain-specific model (CPath-Omni, 72.4% Tiny / 72.2% All).
+While CPathAgent shares patch-level training data with CPath-Omni, its superior performance stems from the third-stage training on high-quality, multi-scale data, which strengthens its analytical capabilities and lays the foundation for agent-based diagnostic workflow.
 CPathAgent even outperforms the expert-annotated baselines in several subsets, suggesting that its general patch-level understanding is already quite promising.
-This may reflect limitations in human annotations, as individual experts often specialize in narrow domains and may struggle with unfamiliar cases in a diverse dataset like PathMMU.
-While CPathAgent benefits from broad training and shows potential as a general-purpose pathology model, patch-level tasks remain relatively simple—real clinical diagnosis often demands more complex reasoning across larger spatial contexts.
+This may also reflect limitations in human annotations, as individual experts often specialize in narrow domains and may struggle with unfamiliar cases in a diverse dataset like PathMMU.
+While CPathAgent benefits from broad training and shows potential as a general-purpose pathology model, patch-level tasks remain relatively simple compared to more complex reasoning across larger spatial contexts.
 Table 1: Overall results of models on the PathMMU test set. The best-performing LMM in each subset for general and pathology LMMs is in-bold, and the second-best performing LMM is underlined.
@@
 4.2 Huge Region Understanding Evaluation
-We evaluate CPathAgent’s ability to autonomously navigate and reason over huge, high-resolution pathology regions using our PathMMU-HR² benchmark, which includes 1,688 expert-validated VQA pairs requiring multi-scale analysis.
-This evaluation is designed to assess whether models can mimic expert-like behaviors—such as zooming in and out, adjusting magnification, and selectively focusing on diagnostically relevant areas.
-Similar to Section 4.1, we involve the same set of models, but one difference is that we apply the prompt used by CPathAgent, enabling stronger agent-capable models like Gemini-2.5-Pro and GPT-4.1-mini to directly mimic the pathologist’s zoom-in/zoom-out movement reasoning process to assess whether this can further enhance their capabilities.
-In addition, we also invited a pathology expert to evaluate both the diagnostic description generation tasks and VQA tasks for 40 regions, including whether their navigation path planning and multi-scale multi-view reasoning aligns with clinical logic.
-We ran the model 8 times to calculate the pass@k.
-Results: CPathAgent demonstrates superior performance on huge, high-resolution pathology regions across diverse cancer types.
-As shown in Table 2, CPathAgent achieves an impressive 88.6% accuracy on the PathMMU-HR² benchmark, substantially outperforming all other models.
-Despite having a more limited single-view capability (1008×1008 resolution) compared to larger models like Gemini-2.5-Pro (3072×3072 resolution), CPathAgent effectively overcomes this limitation by mimicking pathologists’ diagnostic workflow, strategically navigating, zooming in on relevant regions, and integrating observations across multiple views to form comprehensive diagnostic reasoning.
-We provide examples of CPathAgent’s planning and navigation process in Appendix D.
+We evaluate CPathAgent’s ability to autonomously navigate and reason over huge, high-resolution regions using PathMMU-HR², which includes 1,668 expert-validated VQA pairs requiring multi-scale analysis.
+We benchmark the same models as in Section 4.1.
+Additionally, we apply CPathAgent’s prompt (from global screening to multi-scale reasoning, detailed in Section 3.1) to stronger agent-capable models like Gemini-2.5-Pro and GPT-4.1-mini, enabling them to mimic pathologists’ workflows as CPathAgent does and assess whether this approach yields performance gains.
+Results: CPathAgent demonstrates superior performance on large, high-resolution regions across diverse cancer types.
+As shown in Table 2, CPathAgent achieves 88.6% on the PathMMU-HR², substantially outperforming both general-purpose models (Gemini-2.5-Pro by 15.4%) and pathology-specific models (CPath-Omni by 16.9%).
+Despite operating at lower single-view resolution (1008×1008) compared to closed-source models like Gemini-2.5-Pro (3072×3072), CPathAgent compensates through emulating pathologists’ diagnostic workflow through strategic navigation and multi-view reasoning over huge regions.
+Notably, while CPathAgent and CPathOmni share patch-level training data from stages 1-2 (Section 3.4), they differ significantly in their stage 3 training, where CPathAgent incorporates agent-based navigation and reasoning for WSI and large region analysis.
+The substantial performance gain validates the effectiveness of our CPathAgent-Instruct dataset and agent training approach.
+Representative examples are provided in Appendix E.
 Agent-based approaches consistently enhance model performance for pathology analysis.
-As shown in Table 2, the results clearly demonstrate that incorporating agent-like navigation capabilities improves performance across models.
-Overall, models equipped with CPathAgent’s prompting strategy show an average performance improvement of 3.2 percentage points, with Gemini-2.5-Pro improving from 73.2% to 76.4% and GPT-4.1-mini from 60.1% to 62.3% on the full benchmark.
+As shown in Table 2, incorporating agent-based approach improves performance across models.
+Specifically, with CPathAgent’s prompting strategy, Gemini-2.5-Pro improves from 73.2% to 76.4% and GPT-4.1-mini from 60.1% to 62.3%.
 The enhancement is particularly pronounced in complex cancer types like KICH, where Gemini-2.5-Pro shows a dramatic improvement from 56.5% to 69.3% (+12.8%).
-This consistent pattern across models and cancer types confirms that the ability to dynamically explore and reason over pathology images in a manner similar to expert pathologists is crucial for accurate diagnosis, especially for cases requiring integration of findings across multiple scales and regions.
-These promising results suggest that empowering vision models with expert-like navigation strategies could be a key approach for advancing AI performance in complex medical imaging tasks beyond pathology.
-Pathologist evaluation confirms that the navigation paths and diagnostic reasoning generated for both tasks meet clinical expectations.
-As shown in 5, VQA-based approaches achieve higher satisfaction rates than general description tasks.
-We attribute this to VQA’s focused attention on target features, enabling better identification of key elements, while general descriptions require additional judgment regarding which content should be highlighted first.
-Furthermore, as shown in Figure 5, the pass@k metric demonstrates an upward trend as the number of model runs increases, indicating the model’s potential to achieve improved results through multiple generations.
+This confirms that dynamically navigating and reasoning over huge regions that are similar to expert pathologists is crucial for accurate diagnosis, especially for cases requiring integration of findings across multiple scales and regions.
+These results suggest that empowering LMMs with expert-like navigation and reasoning strategies is a promising direction for advancing AI in pathology tasks.
 Table 2: Performance comparison of huge region classification across multiple cancer types using two approaches: General LMMs, and Agent-based Methods.
@@
 PathGen-LLaVA 62.0 77.6 67.4 58.2 74.2 48.4 67.1 84.4 48.1 67.2 67.2
+CPath-Omni 72.6 77.6 71.3 64.1 67.5 59.7 74.3 71.5 76.6 78.8 71.7
 Agent-based Approach
@@
 CPathAgent (Ours) 87.0 88.5 87.8 87.9 92.9 78.9 90.7 89.0 90.7 93.0 88.6
-Figure 5: Success rates (pass@k) across 8 generations, evaluated by pathologists across 40 regions.
-Results show the percentage of pathologist satisfaction with navigation paths and reasoning for General Description Generation (left) and VQA (right) as the number of generations increases.
 4.3 WSI Classification
-We evaluate CPathAgent’s performance on Whole Slide Image (WSI) classification across six diverse datasets spanning multiple cancer types and diagnostic tasks: TCGA-BRCA (breast cancer subtyping), TCGA-NSCLC (lung cancer subtyping), TCGA-RCC (renal cell carcinoma subtyping), TCGA-ESCA (esophageal cancer subtyping), TCGA-BLCA (bladder cancer subtyping), and TCGA-THCA (thyroid cancer subtyping).
-We benchmark CPathAgent against traditional MIL approaches, including ABMIL and DSMIL, as well as both general-purpose and pathology-specific multimodal foundation models.
-Since CPathAgent generates detailed diagnostic descriptions rather than simple classification labels, we implement a two-stage evaluation process to fairly compare with existing benchmarks.
-First, CPathAgent employs our agent-based approach—scanning at low magnification to identify suspicious regions, then strategically analyzing specific areas to generate comprehensive diagnoses for all significant regions within each WSI.
-These detailed descriptions capture the complex pathological findings but must be translated to match the predefined classification schemes of standard benchmarks.
-Therefore, we provide these comprehensive descriptions to Gemini-2.5-Pro for classification into the specific labels required by each dataset.
-The more accurate and detailed CPathAgent’s descriptions, the more precise the resulting WSI classifications.
-Results: CPathAgent achieves competitive performance on whole slide image classification tasks despite using a fundamentally different approach.
-As shown in Table 3, CPathAgent achieves an average accuracy of 82.8% across six different cancer classification tasks, outperforming traditional MIL approaches (ABMIL: 79.9%, DSMIL: 76.8%) and substantially surpassing general-purpose LMMs like Gemini-2.5-Pro (72.1%).
-CPathAgent demonstrates particularly strong performance on TCGA-BRCA (88.5%) and TCGA-ESCA (97.1%), where it matches or exceeds specialized MIL methods.
-While the overall margin over MIL approaches is modest (+2.9% vs.
-ABMIL), this represents a significant paradigm shift—CPathAgent achieves these results through interpretable, expert-like diagnostic reasoning rather than through opaque feature aggregation.
-The agent-based approach demonstrates substantial headroom for further improvement.
-The performance gap between CPathAgent (82.8%) and the theoretical upper bound (91.7%) suggests significant potential for further optimization.
-This upper bound represents the performance achieved by directly using our generated expert-quality data for whole slide image classification, highlighting both the quality of our training data and the considerable margin still available for improvement.
-Despite using high-quality training data, CPathAgent still shows a nearly 9% gap to the overall upper bound, particularly in challenging datasets like TCGA-BLCA and TCGA-THCA.
-This suggests that enhancing agent navigation strategies and reasoning capabilities could yield substantial gains.
+We evaluate CPathAgent’s performance on WSI classification tasks across six diverse datasets spanning multiple cancer types and diagnostic tasks: TCGA-BRCA (breast cancer subtyping), TCGA-NSCLC (lung cancer subtyping), TCGA-RCC (renal cell carcinoma subtyping), TCGA-ESCA (esophageal cancer subtyping), TCGA-BLCA (bladder cancer subtyping), and TCGA-THCA (thyroid cancer subtyping).
+We benchmark CPathAgent against traditional MIL approaches, including ABMIL [4] and DSMIL [5], as well as both general-purpose and pathology-specific LMMs.
+Since CPathAgent generates detailed diagnostic descriptions rather than dataset-specific classification labels like MIL methods, we implement a two-stage evaluation process for fair comparison with WSI benchmarks.
+In the first stage, CPathAgent employs our agent-based approach to identify suspicious regions and generate comprehensive diagnostic summaries of significant findings within each WSI.
+While these detailed descriptions effectively capture complex pathological characteristics, they must be mapped to the predefined classification schemes used in standard benchmarks.
+Therefore, in the second stage, we provide these descriptions to Gemini-2.5-Pro for conversion into the specific labels required by each dataset.
+Consequently, more accurate and detailed diagnostic descriptions from CPathAgent lead to more precise WSI-level classifications.
+For other LMM-based approaches, following CPath-Omni [10], we directly prompt the models to generate descriptions for all WSI regions.
+These descriptions are then similarly provided to Gemini-2.5-Pro to map them to the corresponding classification labels, ensuring fair evaluation across all methods.
+Results: CPathAgent achieves competitive performance on WSI classification using a fundamentally different agent-based approach.
+As shown in Table 3, CPathAgent achieves an average accuracy of 82.8% across six different cancer classification tasks, outperforming traditional MIL approaches (ABMIL: 79.9%, DSMIL: 76.8%) and substantially surpassing both general-purpose LMMs like Gemini-2.5-Pro (72.1%) and previous pathology-specific SOTA CPath-Omni (77.4%).
+CPathAgent shows particularly strong performance on TCGA-BRCA (88.5%) and TCGA-ESCA (97.1%), where it matches or exceeds specialized MIL methods.
+Although the overall margin over MIL approaches is modest (+2.9% vs.
+ABMIL), this represents a significant paradigm shift: CPathAgent achieves these results through interpretable, expert-like diagnostic reasoning rather than opaque feature aggregation.
+Upper bound results validate synthetic data quality and reveal substantial room for improvement.
+To assess data quality, we establish a theoretical upper bound by directly using our generated agent training data for WSI classification (Table 3).
+The near-perfect upper bound accuracy across most datasets (TCGA-ESCA: 100%, TCGA-NSCLC: 97.9%, TCGA-RCC: 96.5%, TCGA-BRCA: 97.0%) confirms that our synthetic navigation trajectories and diagnostic reasoning are clinically accurate and diagnostically complete.
+While CPathAgent achieves significant advances with 82.8% overall accuracy, the 9% gap to the upper bound (91.7%) indicates considerable room for further improvement, especially on challenging datasets like TCGA-BLCA and TCGA-THCA.
+As an agent-based exploration for pathology, CPathAgent demonstrates substantial potential for advancement.
 Table 3: Overall results of WSI classification tasks.
-The best-performing model in each subset for general and pathology domain is in-bold, and the second-best performing model is underlined.
-Balanced accuracy (%) is employed.
+The best-performing model in each subset is in-bold, and the second-best performing model is underlined.
+CPath-Omni* indicates a retrained version using CPathAgent’s train/test split to ensure fair comparison, as the original training data differed.
+Balanced accuracy (%) is reported.
 TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
@@
 DSMIL 84.7 87.2 88.9 89.9 52.2 57.8 76.8
-General Large Multimodal Models
+Large Multimodal Models
 GPT-4.1-mini-2025-04-14 52.8 65.2 52.8 85.3 57.0 54.3 59.6
@@
 Gemini-2.5-Pro-Preview-03-25 72.4 89.2 69.2 97.1 59.8 44.9 72.1
+Quilt-LLaVA 55.3 54.0 53.3 58.6 61.9 45.8 54.8
+PathGen-LLaVA 66.5 53.5 59.0 78.6 54.7 49.2 60.2
+CPath-Omni* 78.6 86.7 91.3 89.9 65.5 52.6 77.4
 Agent-based Approach
 CPathAgent (Ours) 88.5 90.8 94.6 97.1 62.7 63.2 82.8
+Upper Bound 97.0 97.9 96.5 100.0 77.8 81.0 91.7
+Figure 4: OOD results for WSI (left) and region-level (right) classification tasks. Representative models from WSI and region-level approaches are selected for comparison.
+4.4 Out of Distribution Evaluation
+Although our model achieves strong diagnostic performance across multiple TCGA cancer types, we further evaluate its generalization capability on out-of-distribution (OOD) datasets.
+Specifically, we assess CPathAgent on two tasks: (1) binary WSI classification for distinguishing LUAD and LSCC in the CPTAC-Lung dataset [59; 60], and (2) three-class classification (benign tumors, atypical tumors, malignant tumors) on huge image regions (up to 10K $\times$ 10K pixels) in the BRACS dataset [61].
+We benchmark our method against representative WSI-level (PRISM [9], CPath-Omni [10], TITAN [28]) and region-level models (GPT-4.1, Gemini-2.5-Pro, CPath-Omni [10]).
 […]
+While current OOD performance may not yet meet clinical deployment standards, these results demonstrate meaningful generalization beyond TCGA with remarkable data efficiency, suggesting that scaling to larger datasets represents a promising direction for further improvement.
 5 Conclusion
-In this study, we introduced CPathAgent, an agent-based approach for computational pathology that emulates pathologists’ diagnostic reasoning through strategic examination of high-resolution pathology images.
-By meticulously replicating critical diagnostic processes—such as zooming in to inspect cellular details, zooming out to assess tissue architecture, and systematically shifting focus between regions of interest, CPathAgent achieves both enhanced interpretability and state-of-the-art performance across diverse pathology tasks.
-Our comprehensive evaluation across three levels of image scale (patch-level understanding, large regional assessment, and whole slide image analysis) demonstrates that this pathology-mimicking approach, which closely mirrors the trained eye of experienced pathologists navigating tissue specimens, effectively bridges the crucial gap between human diagnostic expertise and computational analysis.
+In this study, we present CPathAgent, an agent-based framework for computational pathology that emulates pathologists’ diagnostic reasoning through strategic examination of high-resolution pathology images.
+Through a systematic three-stage process of global screening, navigation planning, and multi-scale sequence reasoning, CPathAgent faithfully replicates the critical diagnostic workflow by dynamically zooming in and out while systematically shifting focus across regions of interest, transparently documenting not only what features are observed but also why specific regions warrant closer examination and how evidence accumulates across viewing scales, thereby achieving both superior interpretability and state-of-the-art performance across diverse pathology tasks.
+Our comprehensive evaluation spanning three scales of image analysis (patch-level understanding, large regional assessment, and whole slide image analysis) demonstrates that this agent-based paradigm, which closely mirrors how experienced pathologists navigate and scrutinize tissue specimens, effectively bridges the crucial gap between human diagnostic expertise and AI-assisted analysis.
+6 Acknowledgements
+This study was partially supported by "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant 2025SDXHDX0003), the National Natural Science Foundation of China (Grant No.62506306), foundation of Muyuan Laboratory (Program ID: 14106022401,14106022402), and the Westlake Education Foundation.
 Supplementary Material
-[…]
 Appendix A Additional Details of Proposed Dataset
@@
 A.1 Details of CPathAgent-Instruct
-Figure 6: Overview of the VQA-oriented dataset generation process: from multi-scale description to question generation, question filtering, question-oriented navigation path generation, and VQA-related visual reasoning.
-Image Sources: CPathAgent derives its WSI report data from HistGen [57], while the corresponding whole slide image data is sourced from The Cancer Genome Atlas (TCGA).
-TCGA provides a comprehensive collection of pathology whole-slide images contributed by various participating institutions, encompassing a diverse range of distinct tissue and cancer types.
+Figure A1: Overview of the VQA-oriented dataset generation process: from multi-scale description to question generation, question filtering, question-oriented navigation path generation, and VQA-related visual reasoning.
+Data Sources: CPathAgent derives its WSI report data from HistGen [49], while the corresponding WSI data is sourced from The Cancer Genome Atlas (TCGA).
+TCGA provides a comprehensive collection of pathology WSIs contributed by various participating institutions, encompassing a diverse range of distinct tissue and cancer types.
 This extensive repository ensures diverse coverage of pathological conditions and features essential for dataset and model development.
@@
 We provide the following additional dataset construction details that could not be included in the main text due to space constraints:
-1) Region Selection Subset Enhancement: For the region selection subset within CPathAgent-Instruct, we face a challenge as this capability was not developed during earlier training stages (as detailed in Section C), and 5,254 WSIs are insufficient for acquiring this skill.
+1) Global Screening Subset Enhancement: For the region selection subset within CPathAgent-Instruct, we face a challenge as this capability was not developed during earlier training stages (stage 1 and stage 2 as detailed in Section 3.4), and 5,254 WSIs are insufficient for acquiring this skill.
 Considering that region selection is manageable for advanced large models like Gemini-2.5-Pro and doesn’t heavily depend on WSI reports, we expand our data by utilizing all 24,429 TCGA overview images (excluding test data).
@@
 At the same time, it introduces natural variations in how regions are selected, which reflects the inherent subjectivity in how different pathologists might evaluate the same slide.
-The specific prompt templates we use can be found in Section D.5.
-2) Navigation Path Diversity Augmentation: Similarly, for the navigation path planning subset, we recognize that pathologists can follow multiple valid paths to reach accurate diagnoses when examining a large region.
-To capture this diversity, we set the generation temperature to 0.8 and prompt Gemini-2.5-Pro 1-3 times per WSI using the approach described in Section 3.2.
-This generates diverse navigation paths for the same region.
+The specific prompt templates we use can be found in Section E.5.
+2) Navigation Path Diversity Augmentation: Similarly, for the navigation path planning subset, we recognize that pathologistsmay follow multiple valid trajectories when examining a large region to reach accurate diagnoses.
+To capture this inherent diversity, we set the generation temperature to 0.8 and prompt Gemini-2.5-Pro randomly 1-3 times per WSI using the approach described in Section 3.2.
+This stochastic prompting process generates diverse navigation paths for identical regions.
 Correspondingly, the multi-scale multi-image reasoning also changes based on the different navigation paths, creating additional variations in our dataset.
-As a result, we substantially expand our dataset while improving the model’s ability to generate diverse, valid outputs.
-3) VQA-oriented Subset Construction and Quality Control: We systematically construct the VQA subset as shown in Figure 6.
+As a result, we substantially expand our dataset while improving the model’s ability to generate diverse, valid trajectories.
+3) VQA-oriented Subset Construction and Quality Control: We systematically construct the VQA subset as shown in Figure A1.
 First, we use 21 multi-scale descriptions from sub-patches at different magnifications (1x, 2x, 4x), which are generated through WSI-report guided prompting (detailed in Section 3.2).
@@
 This ensures our dataset exclusively targets vision-dependent reasoning tasks that cannot be solved through textual cues alone.
-Figure 7: Illustration of the proportional distribution of each subset within the CPathAgent-Instruct dataset.
-Figure 8: Illustration of the distribution of regions across tissue types.
-Figure 9: Illustration of the distribution of regions across tissue types.
-For questions that pass this validation, we generate navigation plans by providing Gemini-2.5-Pro with the question, multi-scale image descriptions, region descriptions, and full region images.
-Similar to our region description reasoning generation, we combine the validated questions with cropped patches from the planned navigation path and region descriptions.
-This prompts Gemini-2.5-Pro to generate comprehensive reasoning that systematically addresses each question—progressing logically from overall tissue architecture to specific diagnostic features, comparing and eliminating answer options through continuous reasoning to arrive at the correct conclusion.
-Through this question-guided navigation planning and visual reasoning approach, we create a dataset that enables CPathAgent to adapt to diverse diagnostic queries, which allows the model to focus on relevant regions and features based on the specific question.
-Proportion of CPathAgent-Instruct subsets Figure 7 illustrates the distribution of different subsets within the CPathAgent-Instruct dataset.
+Figure A2: Illustration of the proportional distribution of each subset within the CPathAgent-Instruct dataset.
+Figure A3: Illustration of the distribution of regions across tissue types.
+Figure A4: Illustration of the distribution of regions across tissue types.
+For questions that pass validation, we proceed to generate navigation plans through a two-stage process.
+First, we provide Gemini-2.5-Pro with the validated question, multi-scale image descriptions, region descriptions, and full region images to generate a navigation path.
+Second, following the approach established in our region description reasoning generation, we combine the validated question with (1) cropped patches extracted along the planned navigation path and (2) corresponding region descriptions.
+This combined input prompts Gemini-2.5-Pro to generate comprehensive reasoning that systematically addresses the question—progressing logically from overall tissue architecture to specific diagnostic features, comparing and eliminating answer options through continuous reasoning to arrive at the correct conclusion.
+This question-guided navigation and reasoning framework yields a dataset that enables CPathAgent to dynamically adapt to diverse diagnostic queries by focusing on task-relevant regions and features based on the specific question at hand.
+Proportion of CPathAgent-Instruct subsets Figure A2 illustrates the distribution of different subsets within the CPathAgent-Instruct dataset.
 Among them, "Viewing Path Planning" and "Multi-Image Understanding and Reasoning" each constitute 32.8% of the total dataset.
@@
 The VQA-oriented subset, though smaller, is sufficient for model adaptation as it shares structural similarities with the main components.
-Statistics of Source Regions: As shown in Figure 8, CPathAgent dataset encompasses 26 distinct tissue types with a total of 78,658 individual regions.
+Statistics of Source Regions: As shown in Figure A3, CPathAgent dataset encompasses 26 distinct tissue types with a total of 78,658 individual regions.
 Breast tissue represents the highest proportion at 13.4% (10,521 samples), while lymphoid tissue appears with 270 samples.
-Statistics of Overview WSIs within region selection subset: As shown in Figure 9, the CPathAgent dataset consists of 26 distinct tissue types with a total of 24,429 WSI overviews.
+Statistics of Overview WSIs within region selection subset: As shown in Figure A4, the CPathAgent dataset consists of 26 distinct tissue types with a total of 24,429 WSI overviews.
 Lung tissue represents the highest proportion at 12.9% (3,150 WSIs), followed by Breast with 3,110 WSIs and Kidney with 2,672 WSIs.
-Figure 10: Visualization of a word cloud derived from general description-based and VQA-oriented reasoning.
-Figure 11: Visualization of a word cloud derived from questions in VQA-oriented subsets in CPathAgent-instruct and PathMMU-HR².
-Word Frequency in Multi-scale Multi-image Reasoning Subsets: We analyze word frequency distributions in our General description-based (left, Figure 10) and VQA-oriented (right, Figure 10) reasoning approaches.
+Figure A5: Visualization of a word cloud derived from general description-based and VQA-oriented reasoning.
+Figure A6: Visualization of a word cloud derived from questions in VQA-oriented subsets in CPathAgent-instruct and PathMMU-HR².
+Word Frequency in Multi-scale Multi-image Reasoning Subsets: We analyze word frequency distributions in our General description-based (left, Figure A5) and VQA-oriented (right, Figure A5) reasoning approaches.
 The visualization reveals distinctive reasoning patterns in pathology diagnosis.
@@
 Both approaches incorporate critical pathological terminology ("stroma," "nuclear," "infiltrate") essential for accurate diagnosis, demonstrating that these reasoning strategies center on key diagnostic elements fundamental to clinical pathology assessment and disease classification.
-Word Frequency of Questions in VQA-oriented Subset: We analyze the question terminology in our VQA-oriented subset, as shown in the left part of Figure 11.
+Word Frequency of Questions in VQA-oriented Subset: We analyze the question terminology in our VQA-oriented subset, as shown in the left part of Figure A6.
 The visualization reveals "diagnosis" as the central focus, surrounded by practical terms like "describe," "integrate," and "feature." Magnification terms ("40x," "10x") indicate how pathologists navigate between different viewing scales, while specific organs ("thyroid," "breast," "lung") appear frequently in questions.
@@
 A.2 Details of PathMMU-HR²
-The construction of PathMMU-HR² is similar to the construction of the VQA-oriented subset in CPathAgen-Instruct, as shown in Figure 6.
+The construction of PathMMU-HR² is similar to the construction of the VQA-oriented subset in CPathAgen-Instruct, as shown in Figure A1.
 However, they differ in that after filtering VQA pairs that can be solved through text-only shortcuts, we engage three pathologists with over 10 years of experience for human validation.
@@
 Questions failing any criterion are deemed invalid.
-Under pathologists’ evaluation, we identify 1,688 valid questions from the initially annotated 2,200 questions, with the annotation platform interface shown in Figure 12.
-We also visualize the question part of PathMMU-HR² with a word cloud in Figure 11.
+Under pathologists’ evaluation, we identify 1,668 valid questions from the initially annotated 2,200 questions, with the annotation platform interface shown in Figure A7.
+We also visualize the question part of PathMMU-HR² with a word cloud in Figure A6.
 As shown, the most prominent terms—such as "diagnose," "integrate," "accurately," and "across"—highlight that many queries require multi-scale analysis and complex reasoning.
@@
 This multi-scale integration is essential for correctly answering questions that demand the model to synthesize information from subcellular details to tissue architecture, evaluating both cytological features and architectural patterns simultaneously to arrive at accurate answers.
-Figure 12: Example of the PathMMU-HR2 annotation interface.
+Figure A7: Example of the PathMMU-HR2 annotation interface.
 Appendix B Additional Experiments and Discussion
@@
 Results: CPathAgent significantly outperforms both PathGen-LLaVA and Quilt-LLaVA across all evaluation metrics.
-As shown in Figure 13, most notably, CPathAgent achieves a perfect 100% win rate against Quilt-LLaVA across all three evaluators.
-Against PathGen-LLaVA, CPathAgent maintains consistently strong performance with over 95% win rates across all evaluators.
+As shown in Figure A8, most notably, CPathAgent achieves a perfect 100% win rate against Quilt-LLaVA across all three evaluators.
+Against PathGen-LLaVA, CPathAgent maintains consistently strong performance with over 96% win rates across all evaluators.
 The evaluation consistency between human and LMM evaluators is remarkably high.
@@
 General-purpose LMMs, may favor CPathAgent’s comprehensive reasoning, while the pathologist evaluator places equal weight on both reasoning quality and diagnostic accuracy, resulting in a more reliable assessment that recognizes PathGen-LLaVA’s clinical precision in some cases.
-Figure 13: Comparison of CPathAgent with different pathology-specific LMMs for description/diagnosis generation on 100 randomly sampled cases, evaluated by Gemini-2.5-Pro, GPT-4.1, and expert pathologists.
-B.2 Comparison of WSI Classification Performance Among Pathology-specific LMMs
-We compare CPathAgent with existing pathology-specific LMMs on WSI classification tasks, following the same approach as used for general LMMs detailed in Section 4.3.
-Results: Both PathGen-LLaVA and Quilt-LLaVA significantly underperform on WSI classification, achieving only 60.2% and 54.8% overall balanced accuracy, respectively, as shown in Table 4.
-This performance gap can be attributed to two key limitations of these models.
-First, PathGen-LLaVA and Quilt-LLaVA are specifically designed and trained for patch-level analysis using small tissue regions.
-Consequently, they lack the ability to capture high-level, contextual diagnostic information that is crucial for WSI analysis.
-Second, these models do not possess the capability to navigate and zoom within large-scale regions.
-When large regions are resized to fit their input requirements, critical diagnostic details become virtually indiscernible, severely compromising their performance on WSI classification tasks.
-Table 4: Overall results of WSI classification tasks.
-The best-performing model in each subset for general and pathology domain is in-bold, and the model with the second-best performance is underlined.
-Balanced accuracy (%) is employed.
+Figure A8: Comparison of CPathAgent with different pathology-specific LMMs for description/diagnosis generation on 100 randomly sampled cases, evaluated by Gemini-2.5-Pro, GPT-4.1, and expert pathologists.
+B.2 Pathologist evaluation results
+We invited a pathology expert to evaluate both diagnostic description generation and VQA tasks across 40 regions, assessing whether the navigation path planning and multi-scale multi-view reasoning align with clinical logic.
+The model was run 8 times to calculate pass@k metrics.
+Results: Pathologist evaluation confirms that both navigation paths and diagnostic reasoning meet clinical expectations.
+As shown in Figure A9, the model demonstrates strong performance with both tasks.
+For navigation path evaluation, CPathAgent’s initial satisfaction rates are acceptable but modest, starting at 70% for general description generation and 77.5% for VQA.
+However, the pass@k metric demonstrates promising improvement when increasing model runs, with both tasks ultimately exceeding 85% satisfaction.
+For reasoning evaluation, satisfaction rates start higher at 82.5% for general description generation and 87.5% for VQA, and show continued improvement, reaching 90% for general description generation and 92.5% for VQA at higher k values.
+Across both tasks, the consistent upward trend in pass@k metrics suggests that the model can reliably generate clinically sound navigation strategies and reasoning when given multiple attempts.
+Additionally, VQA-based approaches consistently outperform general description tasks across both navigation planning and reasoning.
+We attribute this to VQA’s focused attention on target features that enables better identification of key elements, whereas general descriptions require additional judgment about content prioritization.
+Figure A9: Success rates (pass@k) across 8 generations, evaluated by pathologists across 40 regions.
+Results show the percentage of pathologist satisfaction with navigation paths and reasoning for General Description Generation (left) and VQA (right) as the number of generations increases.
+B.3 Additional Results on PathMMU-Pro
+While CPathAgent demonstrates strong performance on PathMMU, achieving results that exceed the reported expert baselines, we acknowledge that the observed gap between our model and expert performance can be attributed to two primary factors inherent to the benchmark design.
+First, the expert annotations in PathMMU were conducted by pathologists with specialized expertise in one or two specific domains, whereas the dataset encompasses a broad spectrum of tissue types and pathological conditions across multiple subspecialties.
+Consequently, no individual expert possessed comprehensive knowledge across all domains represented in the evaluation set.
+Second, as discovered by the PathMMU, large multimodal models may exploit text-based reasoning shortcuts, leveraging strong language priors and contextual cues from question formulations that are not readily accessible to human experts performing traditional diagnostic tasks.
+To address the concern regarding text-based shortcuts and provide a more rigorous assessment of our model’s genuine pathological reasoning capabilities, we conducted additional evaluation on PathMMU-Pro [63], a refined version of the benchmark specifically designed with more carefully constructed questions that minimize textual biases.
+Results: On PathMMU-Pro, CPathAgent no longer surpasses expert performance but maintains a substantial lead over the previous SOTA CPath-Omni.
+As shown in Table A4, expert performance now leads with an average accuracy of 69.4%, surpassing CPathAgent’s results.
+This shift substantiates that the more rigorous question design in PathMMU-Pro effectively mitigates the text-based guessing capabilities that LMMs could previously exploit.
+Nevertheless, CPathAgent demonstrates consistent advantages over CPath-Omni, achieving 66.3% versus 62.1% on the Test-Tiny subset and 65.3% versus 61.8% on the Test-All subset, demonstrating that CPathAgent possesses stronger multimodal understanding capabilities for pathology.
+Table A4: Overall results of models on the PathMMU-Pro. The best-performing model in each subset is in-bold.
+Test Overall PubMed SocialPath EduContent
+Tiny ALL Tiny ALL Tiny All Tiny All
+(430) (4755) (147) (1865) (136) (1081) (147) (1809)
+Expert performance 69.4 - 71.2 - 70.1 - 66.9 -
+CPath-Omni 62.1 61.8 60.5 60.8 58.8 62.1 66.7 63.0
+CPathAgent (Ours) 66.3 65.3 64.6 65.8 65.4 63.6 68.7 66.2
+B.4 Ablation Study on Agent Components
+To assess the contribution of each agent component, we conduct systematic ablation studies by isolating individual module impacts.
+We evaluate two critical components.
+For global screening, we replace the learned region selection module with random selection of 70% of WSI regions.
+For navigation planning and multi-scale reasoning, we ablate the dynamic navigation and multi-view analysis by directly generating descriptions from entire regions at once, followed by WSI classification based on these single-view descriptions.
+Results: Both agent components contribute meaningfully to performance, with navigation and multi-scale reasoning being particularly critical for diagnostic accuracy.
+As shown in Table A5, removing strategic global screening leads to a 2.5% performance drop (from 82.8% to 80.3%), demonstrating that learned region selection effectively identifies diagnostically informative areas and mimics pathologists’ selective attention during screening.
+The impact of removing navigation and multi-scale reasoning is substantially more pronounced, with performance degrading by 7.8% (from 82.8% to 75.0%).
+This significant drop reveals two key limitations of direct WSI region processing: (1) directly processing high-resolution regions (up to $\times16000$ pixels) necessitates aggressive downsampling to fit model input constraints, causing critical diagnostic details to become indiscernible; (2) the absence of pathologist-like reasoning that strategically examines multiple fields of view at varying magnifications fundamentally limits the model’s diagnostic capability.
+These results confirm that CPathAgent’s effectiveness stems from the integration of all components, with navigation planning and multi-scale reasoning being particularly essential for achieving high diagnostic accuracy on WSI tasks.
+Table A5: Results of ablation study on agent components for WSI classification. Balanced accuracy (%) is reported.
+Model TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
+Full CPathAgent 88.5 90.8 94.6 97.1 62.7 63.2 82.8
+w/o Gloabal Screening 84.7 88.7 94.6 94.1 60.8 58.9 80.3
+w/o Navigation & Reasoning 80.0 83.6 88.7 94.1 52.2 51.3 75.0
+Table A6: Results of ablation study on applying Gemini-2.5-Pro as an agent for WSI tasks. Balanced accuracy (%) is reported
 TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
 General Large Multimodal Models
-GPT-4.1-mini-2025-04-14 52.8 65.2 52.8 85.3 57.0 54.3 59.6
-GPT-4.1-2025-04-14 61.0 65.8 53.8 85.3 61.3 58.9 64.3
-Gemini-2.5-Flash-Preview-04-17 56.7 84.5 57.9 93.8 61.8 55.4 68.4
 Gemini-2.5-Pro-Preview-03-25 72.4 89.2 69.2 97.1 59.8 44.9 72.1
-Pathology-specific Large Multimodal Models
-Quilt-LLaVA 55.3 54.0 53.3 58.6 61.9 45.8 54.8
-PathGen-LLaVA 66.5 53.5 59.0 78.6 54.7 49.2 60.2
 Agent-based Approach
@@
 CPathAgent (Ours) 88.5 90.8 94.6 97.1 62.7 63.2 82.8
-[…]
-B.3 Ablation Study on Applying Gemini-2.5-Pro as an Agent for WSI Tasks
+B.5 Ablation Study on Applying Gemini-2.5-Pro as an Agent for WSI Tasks
 We also apply CPathAgent’s inference prompts to directly prompt Gemini-2.5-Pro, enabling it to perform agent-based reasoning and diagnosis similar to CPathAgent.
-Results: Interestingly, while this approach yields modest improvements on TCGA-BLCA and TCGA-THCA datasets, the overall performance decreases by 6.7% (from 72.1% to 65.4%), as shown in the "Agent-based Approach" part of Table 4.
+Results: Interestingly, while this approach yields modest improvements on TCGA-BLCA and TCGA-THCA datasets, the overall performance decreases by 6.7% (from 72.1% to 65.4%), as shown in the "Agent-based Approach" part of Table A6.
 This contrasts with the findings in Section 4.2 of the main paper, where applying agent-based approaches to advanced LMMs on large regions VQA demonstrated performance gains.
@@
 These targeted descriptions enable us to prompt Gemini-2.5-Pro to capture multi-scale morphological features across different viewing perspectives.
-By understanding which anatomical locations contain specific pathological features, we can further prompt Gemini-2.5-Pro to generate higher-quality navigation paths with enhanced spatial awareness, which in turn enables more precise and accurate multi-scale, multi-view reasoning generation.
+By understanding which locations contain specific pathological features, we can further prompt Gemini-2.5-Pro to generate higher-quality navigation paths with enhanced spatial awareness, which in turn enables more precise and accurate multi-scale, multi-view reasoning generation.
 This creates a hierarchical sequence where each stage builds upon the quality of previous stages, resulting in cascading improvements throughout the entire data generation pipeline.
@@
 Models with the best performance on the validation set are selected for evaluation on the test set.
-C.1.2 Details for WSI preprocessing
+C.1.2 Details for MIL preprocessing
 Our WSI preprocessing pipeline follows the approach established by CPath-Omni [10] to implement a multi-scale MIL-based methodology.
@@
 The MIL framework commonly used for WSI classification includes three learnable components: (1) a fully-connected layer to reduce the dimensionality of features to 256, (2) an attention network to aggregate and transform the instance features, and (3) a final fully-connected layer for making predictions.
-We experiment with ABMIL [61] and DSMIL [5], both of which share the same fully-connected layers for reducing feature dimensionality and making predictions.
+We experiment with ABMIL [4] and DSMIL [5], both of which use the same fully-connected layers for reducing feature dimensionality and making predictions.
 For the attention network, ABMIL uses a gated attention mechanism, while DSMIL introduces a dual-stream architecture.
-C.2 Hardware and Training Cost of CPathAgent
+C.2 Details for Patch Level Benchmark
+PathMMU [48]: An expert-validated pathology benchmark comprising 33,428 multimodal multiple-choice questions and 24,067 images from diverse sources including PubMed, pathology atlases, expert social media posts, and educational content.
+Constructed using GPT-4V with rigorous validation by seven pathologists, it covers multiple organ systems and pathology subjects for expert-level evaluation of LMMs’ pathology understanding and reasoning capabilities.
+PathMMU-Pro [63]: Building upon PathMMU, PathMMU-Pro introduces more strict assessment to address a critical limitation in multimodal evaluation: the potential for models to solve questions using text-only reasoning without truly leveraging visual information.
+PathMMU-Pro employs a systematic two-step refinement process: (1) training specialized text-only "question-guessing" models that receive only textual inputs to identify and filter out questions answerable without examining pathology images, and (2) enhancing remaining questions by generating more confusing options that are textually similar but require careful visual examination to distinguish, thereby preventing models from guessing answers through superficial text patterns.
+This enhanced benchmark provides a more strict assessment of models’ intrinsic visual reasoning capabilities in pathology.
+C.3 Hardware and Training Cost of CPathAgent
 We employ 8 NVIDIA H800-80G GPUs for CPathAgent training, with computational requirements of 9 hours for Stage 1 (multimodal alignment), 25 hours for Stage 2 (pathology task training), and 39 hours for Stage 3 (agent capability development).
-C.3 Training Hyperparameters
+C.4 Training Hyperparameters
 For Stages 1 and 2 of CPathAgent training, we follow the hyperparameter configuration of CPath-Omni.
@@
 We implement a cosine learning rate scheduler with a base learning rate of 1e-5 and a warmup ratio of 0.03, while fine-tuning the vision tower with a reduced learning rate of 2e-6.
-Appendix D Examples and Prompts
-D.1 Examples of Region Descriptions Extracted from WSI Report Pairs
-As shown in Figure 14, we extract relevant region descriptions from WSI reports by identifying specific diagnostic findings in targeted areas.
+Appendix D Limitations
+While CPathAgent successfully demonstrates the feasibility of mimicking pathologists’ diagnostic reasoning pathways, several limitations merit future improvement:
+1.
+Synthetic diagnostic pathways.
+Currently, the diagnostic viewing paths employed by CPathAgent are synthetically generated rather than derived from real pathologist workflows.
+Although our synthetic paths are designed based on established diagnostic principles and demonstrate effectiveness in our experiments, they may not fully capture the nuanced decision-making processes and adaptive strategies employed by experienced pathologists in clinical practice.
+Future work should incorporate authentic diagnostic paths collected from pathologists through eye-tracking studies or explicit annotation of their reasoning sequences.
+2.
+Limited task scope.
+CPathAgent currently focuses exclusively on diagnostic classification tasks.
+While this represents a critical clinical application, computational pathology encompasses a broader spectrum of tasks including prognostic prediction and biomarker detection.
+Extending CPathAgent to these domains is theoretically feasible and would significantly expand its clinical utility.
+3.
+Training data scale.
+Our model was trained on 5,254 WSIs, which, while sufficient to demonstrate proof-of-concept, remains considerably smaller than datasets used by state-of-the-art foundation models such as PRISM (587,196 WSIs) and TITAN (335,645 WSIs).
+This disparity in scale may limit the model’s generalization capabilities and its ability to capture rare morphological patterns.
+Scaling up the training data represents a clear path toward improved performance and robustness.
+Appendix E Examples and Prompts
+E.1 Examples of Region Descriptions Extracted from WSI Report Pairs
+As shown in Figure A10, we extract relevant region descriptions from WSI reports by identifying specific diagnostic findings in targeted areas.
 This process systematically parses pathological reports to highlight key diagnostic information related to the given region, addressing the limitation in TCGA datasets where WSI reports exist but region-specific descriptions are missing.
-D.2 Examples of PathMMU-HR² Samples
-Figure 15 shows the examples of PathMMU-HR² dataset.
-These VQA samples demonstrate exceptional quality through the requirement for multi-scale, multi-view reasoning to arrive at accurate diagnoses.
+E.2 Examples of PathMMU-HR² Samples
+Figure A11 shows the examples of PathMMU-HR² dataset.
+These VQA samples require multi-scale, multi-view reasoning to arrive at accurate diagnoses.
 As illustrated in the examples, the correct identification of conditions such as fibromatosis (desmoid tumor) and well-demarcated clear cell tumors necessitates careful examination of architectural patterns and cellular features across different magnification levels.
@@
 This multi-scale analytical approach reflects real-world pathological practice, where definitive diagnoses depend on the integration of macro-architectural features with microscopic cellular details.
-D.3 Examples of Region Selection Process Based on WSI Overview
-Figure 16 and Figure17 demonstrate the generated region selection results.
+E.3 Examples of Region Selection Process Based on WSI Overview
+Figure A12 and FigureA13 demonstrate the generated region selection results.
 As illustrated, the model segments the WSI overview into distinct regions and identifies which areas require high-magnification examination.
-For example, in Figure 16, the model correctly prioritizes core tumor areas for detailed analysis while also identifying critical tumor interface/periphery regions and margin assessment areas to evaluate potential residual tumor presence.
+For example, in Figure A12, the model correctly prioritizes core tumor areas for detailed analysis while also identifying critical tumor interface/periphery regions and margin assessment areas to evaluate potential residual tumor presence.
 In addition, the model appropriately excludes the benign/background breast tissue areas from high-magnification requirements, as these regions primarily consist of adipose tissue and fibrous stroma that appear benign and uninvolved at low magnification, thus optimizing the diagnostic workflow by focusing computational resources on clinically significant regions.
-D.4 Examples of Navigation Paths and Multi-Scale Reasoning Generated by CPathAgent
-D.4.1 General Navigation Path Planning and Reasoning
-Figures 18 and 19 show how CPathAgent navigates and analyzes pathology images like an expert pathologist.
-For example, as illustrated in Figure 18, CPathAgent begins with a comprehensive overview at low magnification (1.0x), scanning the entire tissue to identify distinct regions and architectural patterns.
+E.4 Examples of Navigation Paths and Multi-Scale Reasoning Generated by CPathAgent
+E.4.1 General Navigation Path Planning and Reasoning
+Figures A14 and A15 show how CPathAgent navigates and analyzes pathology images like an expert pathologist.
+For example, as illustrated in Figure A14, CPathAgent begins with a comprehensive overview at low magnification (1.0x), scanning the entire tissue to identify distinct regions and architectural patterns.
 In this case, it recognizes adipose tissue on the left and dense cellular areas on the right, establishing the overall tissue landscape.
@@
 This systematic, multi-scale approach allows CPathAgent to integrate architectural observations from low magnification with detailed cellular features from high magnification, ultimately reaching a comprehensive diagnosis of invasive ductal carcinoma with appropriate histological grading—mirroring the thorough diagnostic process used by expert pathologists in clinical practice.
-D.4.2 VQA-oriented Navigation Path Planning and Reasoning
-Figure 20 demonstrates CPathAgent’s question-guided reasoning approach for VQA task.
+E.4.2 VQA-oriented Navigation Path Planning and Reasoning
+Figure A16 demonstrates CPathAgent’s question-guided reasoning approach for VQA task.
 When presented with a specific question about "morphological patterns describing the relationship between neoplastic cell populations and stromal elements at their interface," CPathAgent tailors its entire analysis strategy to address this inquiry.
@@
 This question-driven navigation allows CPathAgent to methodically evaluate multiple-choice options, ultimately selecting the answer that best describes the observed "nested clear cells focally transitioning to solid eosinophilic cells with high-grade nuclei and dense lymphocytes." This demonstrates how CPathAgent’s reasoning process is specifically optimized to answer targeted pathological questions through systematic, multi-scale analysis.
-Figure 14: Examples of region description extraction from WSI reports.
-Figure 15: Examples of VQAs from the PathMMU-HR² dataset.
-Figure 16: An example of region selection based on a WSI overview.
-Figure 17: An example of region selection based on a WSI overview.
-Figure 18: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
-Figure 19: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
-Figure 20: An example of CPathAgent’s VQA-oriented navigation path planning and multi-scale, multi-view reasoning process.
-D.5 Prompts
+Figure A10: Examples of region description extraction from WSI reports.
+Figure A11: Examples of VQAs from the PathMMU-HR² dataset.
+Figure A12: An example of region selection based on a WSI overview.
+Figure A13: An example of region selection based on a WSI overview.
+Figure A14: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
+Figure A15: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
+Figure A16: An example of CPathAgent’s VQA-oriented navigation path planning and multi-scale, multi-view reasoning process.
+E.5 Prompts
 This section presents the prompts used in our dataset construction and experiments.
 Gemini-2.5-Pro Prompts for Dataset Construction: We design a series of prompts to guide Gemini-2.5-Pro in constructing our pathologist-like planning and reasoning dataset generation pipeline.
-Figure 21 presents the prompt used to extract region-relevant descriptions from WSI reports, while Figure 22 shows the prompt for region selection using WSI overview images and paired reports.
-For multi-scale image description generation, Figure 23 displays the prompt that generates detailed descriptions for image patches at different magnifications.
-Figure 24 contains the prompt for creating navigation plans using coordinate-annotated region overviews and multi-scale descriptions.
-Finally, Figure 25 shows the prompt that generates pathologist-like diagnostic reasoning based on extracted patches from navigation paths and paired region descriptions.
-For VQA dataset construction, we use several specialized prompts: Figure 26 shows the prompt that instructs the model to generate VQA samples, Figure 27 presents the prompt for predicting question answers from text-only inputs (applied to both Gemini-2.5-Pro and DeepSeek-V3), Figure 28 displays the prompt that guides the model to create question-guided navigation plans, and Figure 29 contains the prompt for generating reasoning-based answers.
+Figure A17 presents the prompt used to extract region-relevant descriptions from WSI reports, while Figure A18 shows the prompt for region selection using WSI overview images and paired reports.
+For multi-scale image description generation, Figure A19 displays the prompt that generates detailed descriptions for image patches at different magnifications.
+Figure A20 contains the prompt for creating navigation plans using coordinate-annotated region overviews and multi-scale descriptions.
+Finally, Figure A21 shows the prompt that generates pathologist-like diagnostic reasoning based on extracted patches from navigation paths and paired region descriptions.
+For VQA dataset construction, we use several specialized prompts: Figure A22 shows the prompt that instructs the model to generate VQA samples, Figure A23 presents the prompt for predicting question answers from text-only inputs (applied to both Gemini-2.5-Pro and DeepSeek-V3), Figure A24 displays the prompt that guides the model to create question-guided navigation plans, and Figure A25 contains the prompt for generating reasoning-based answers.
 CPathAgent Training and Inference Prompts: The prompts for CPathAgent follow a similar design, while removing the additional information utilized by Gemini-2.5-Pro (such as pathology reports), focusing on vision-only pathological analysis that replicates clinical diagnostic processes.
-Figures 30, 31, and 32 present the core prompts used for CPathAgent’s training and inference.
+Figures A26, A27, and A28 present the core prompts used for CPathAgent’s training and inference.
 These prompts guide CPathAgent to perform navigation planning and reasoning using only visual input.
-The VQA-specific prompts shown in Figures 33 and 34 display the prompts that used for CPathAgent’s question-guided navigation and reasoning.
-Figure 21: Prompt for Gemini-2.5-Pro to extract description related to a given region from the corresponding WSI paired pathology report.
-Figure 22: Prompt for Gemini-2.5-Pro to generate region selection results given a WSI overview and its paired pathology report.
-Figure 23: Prompt for Gemini-2.5-Pro to generate descriptions for 1 image at 1X scale, 4 images at 2X scale, and 16 images at 4X scale, given these images and their paired region descriptions.
-Figure 24: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates and 21 multi-scale image descriptions generated in the previous step.
-Figure 25: Prompt for Gemini-2.5-Pro to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path and the overall region description extracted from the previous step.
-Figure 26: Prompt for Gemini-2.5-Pro to generate VQA samples given the region overview and the 21 multi-scale image descriptions generated from the previous step.
-Figure 27: Prompt for Gemini-2.5-Pro to conduct educated guesses given only the text portion of the generated VQA samples.
-Figure 28: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates, 21 multi-scale image descriptions and the question generated in the previous step.
-Figure 29: Prompt for Gemini-2.5-Pro to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path, the overall region description extracted from the previous step, and the target question.
-Figure 30: Prompt for CPathAgent to generate region selection results given a WSI overview.
-Figure 31: Prompt for CPathAgent to generate navigation path planning given the region overview with coordinates.
-Figure 32: Prompt for CPathAgent to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path.
-Figure 33: Prompt for CPathAgent to generate VQA-oriented navigation path planning given the region overview with coordinates and the question.
-Figure 34: Prompt for CPathAgent to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path and the target question.
+The VQA-specific prompts shown in Figures A29 and A30 display the prompts that used for CPathAgent’s question-guided navigation and reasoning.
+Figure A17: Prompt for Gemini-2.5-Pro to extract description related to a given region from the corresponding WSI paired pathology report.
+Figure A18: Prompt for Gemini-2.5-Pro to generate region selection results given a WSI overview and its paired pathology report.
+Figure A19: Prompt for Gemini-2.5-Pro to generate descriptions for 1 image at 1X scale, 4 images at 2X scale, and 16 images at 4X scale, given these images and their paired region descriptions.
+Figure A20: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates and 21 multi-scale image descriptions generated in the previous step.
+Figure A21: Prompt for Gemini-2.5-Pro to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path and the overall region description extracted from the previous step.
+Figure A22: Prompt for Gemini-2.5-Pro to generate VQA samples given the region overview and the 21 multi-scale image descriptions generated from the previous step.
+Figure A23: Prompt for Gemini-2.5-Pro to conduct educated guesses given only the text portion of the generated VQA samples.
+Figure A24: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates, 21 multi-scale image descriptions and the question generated in the previous step.
+Figure A25: Prompt for Gemini-2.5-Pro to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path, the overall region description extracted from the previous step, and the target question.
+Figure A26: Prompt for CPathAgent to generate region selection results given a WSI overview.
+Figure A27: Prompt for CPathAgent to generate navigation path planning given the region overview with coordinates.
+Figure A28: Prompt for CPathAgent to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path.
+Figure A29: Prompt for CPathAgent to generate VQA-oriented navigation path planning given the region overview with coordinates and the question.
+Figure A30: Prompt for CPathAgent to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path and the target question.